### v1

# Weakly-supervised diagnosis identification from Italian discharge letters

Abstract
Objective: Identifying the diagnosis that determines a patient’s hospitalization or Emergency Room admission is crucial for cohort selection and epidemiological analyses. Discharge letters are the only type of data consistently produced across hospitals. Recognizing diseases from discharge letters is a classic document classification problem, typically requiring supervised learning. However, manual annotation of large datasets of discharge letters is uncommon since it is extremely time-consuming. We propose a novel weakly-supervised pipeline to recognize diseases from Italian discharge letters.
Methods: Our Natural Language Processing pipeline is based on a fine-tuned version of the Italian Umberto model. The pipeline extracts diagnosis-related sentences from a subset of letters and applies a two-level clustering using the embeddings generated by the fine-tuned Umberto model. These clusters are summarized and those mapped to the diseases of interest are selected as weak labels. Finally, the same BERT-based model is trained using these weak labels to detect the targeted diseases.
Results: Results on a case study related to the identification of bronchiolitis in a dataset of 33’176 Italian discharge letters from multiple hospitals in the Veneto Region demonstrate the potential of our method, with an AUC of 77.7 % and an F1-Score of 75.1 % on manually annotated labels, improving compared to other non-supervised methods and with a limited loss compared to fully supervised methods. Results are robust to the cluster selection and the identified clusters highlight the potential to recognize a variety of diseases.
Conclusions: This study demonstrates the feasibility of successfully addressing the challenge of diagnosis identification from Italian discharge letters in the absence of labelled data. Our weakly-supervised pipeline showed strong performance and robustness, and its flexibility allows for easy adaptation to various diseases. This approach offers a scalable solution for clinical text classification, significantly reducing the need for manual annotation while maintaining good accuracy.
keywords:
Natural Language Processing , weakly-supervised learning , discharge letters , diagnosis identification
MSC:
68T50
[…]
[inst1]organization=MOX - Modelling and Scientific Computing Lab, Department of Mathematics, Politecnico di Milano,city=Milan, country=Italy
[…]
[inst2]organization=Division of Pediatric Infectious Diseases, Department for Woman and Child Health, University of Padova,city=Padova, country=Italy
[…]
[inst3]organization=Unit of Biostatistics, Epidemiology and Public Health, Department of Statistics and Quantitative Methods, University of Milano-Bicocca,city=Milan, country=Italy
[…]
[inst4]organization=National Centre for Healthcare Research and Pharmacoepidemiology, Department of Statistics and Quantitative Methods, University of Milano-Bicocca,city=Milan, country=Italy
[…]
[inst5]organization=HDS - Health Data Science Centre, Human Technopole,city=Milan, country=Italy
1 Introduction
Electronic Health Records contain a vast and continuously growing amount of patient information, but a significant portion of this data remains unstructured in formats like clinical notes, reports, discharge letters and referrals. Despite containing valuable information, this unstructured data is often underutilized in the downstream analysis due to their complexity [1, 2]. While Natural Language Processing (NLP) has made significant advancements in extracting and classifying information from text [3, 4], applying these techniques to healthcare data presents additional challenges, such as a specialized lexicon, a high prevalence of ambiguous abbreviations, frequent misspellings, and a lack of standard templates across hospitals [5, 6]. Additionally, most NLP tools are developed for high-resource languages, making them less effective for languages like Italian, where linguistic resources are scarce [7].
A critical piece of information in medical documents is the reason for a patient’s hospital admission or discharge, typically a disease diagnosis or a description of the patient’s symptoms. This information is critical for both cohort selection and epidemiological analysis. Although an international code system for diseases and symptoms, the ICD-CM [8], exists, physicians often prefer to provide a textual description of their diagnoses. Textual descriptions are faster to write, more detailed, and are often used because many medical document templates do not include structured fields for disease codes [9, 10]. Even when present, it has been observed that ICD codes are not completely reliable [11, 12, 13]. Consequently, identifying patients admitted for specific diseases or symptoms often requires manual inspection of these texts, which is labour-intensive and impractical at scale.
In Italy, every hospitalization or Emergency Room (ER) admission generates a discharge letter. Depending on the type of examinations and the hospital, more structured data may also be recorded. A discharge letter should always contain information about the diseases that affect the patient, but the complexity of these documents and the lack of standard templates pose a significant challenge for automatic data extraction. To address this issue, this study aims to develop an NLP pipeline capable of extracting structured information from the unstructured text of Italian discharge letters.
Traditional approaches to this problem typically frame it as a supervised classification task, which requires a substantial amount of annotated data to perform well [14]. However, annotating thousands of discharge letters is time-consuming, especially for clinicians who are already overworked. This highlights a significant gap in the field: the need for methods that can effectively utilize large, unannotated datasets without the burden of manual labelling.
While there have been a few studies on NLP applied to Italian healthcare documents, none have focused specifically on discharge letters. Lanera et al. [15, 16] worked on supervised models to identify cases of varicella, using textual data from specific children’s health project dataset. Similarly, Sciannameo et al. [17] used a BERT-based model to predict rehospitalizations for elderly people from the text of medication prescriptions and hospitalization diagnoses. Hammami et al.  [18] classified pathology reports for cancer morphology, developing a rule-based system. Other studies have concentrated on extracting events and entities, such as Viani et al. [19, 20], who analyzed cardiology reports using ontologies and recurrent neural networks.
Significant research on discharge letters has been done in the English language, often using the MIMIC dataset [21, 22] (see [23, 24, 25, 26] for some examples).
Research has also been conducted on discharge letters in other under-resourced languages, such as Bulgarian [27, 28], Dutch [29] and Swedish [30].
All these studies have relied either on large annotated datasets or on the development of complex, problem-specific rule-based systems. In contrast, our approach eliminates the dependency on manually annotated data and offers a more generalizable solution than traditional rule-based systems.
The main contribution of this paper is the first NLP-based pipeline for the weakly-supervised classification of Italian discharge letters, offering a scalable solution to a long-standing challenge in healthcare data management.
The results presented in this paper pertain to a clinically relevant case study aimed at identifying bronchiolitis, one of the most relevant diseases that can determine hospitalizations in young children [31]. An automated model for this purpose can significantly aid in patient cohort selection for research, reducing the manual workload for physicians. The dataset for this case study comprises 33’176 discharge letters related to children from 44 different hospitals in the Veneto region, one of the most populated Italian regions.
The rest of the paper is structured as follows: in Section 2, we discuss the methodology we developed and the data of our case study; in Section 3, we present the results; in Section 4 we discuss them, together with possible future developments, and in Section 5 we summarize the main findings.
Statement of significance
Problem Identifying diagnoses from discharge letters is relevant for both cohort selection and epidemiological analysis but requires large annotated datasets, rarely available in this domain.
What is Already Known Supervised learning methods for diagnosis identification are well-studied, with state-of-the-art methods based on transformers. Unsupervised methods are mainly rule-based algorithms that lack generalizability.
What this Paper Adds This paper introduces a weakly-supervised NLP pipeline that eliminates the need for manual annotations, offering a scalable solution for diagnosis identification. The approach is validated on a case study with a large dataset of Italian discharge letters and demonstrates strong performance and potential generalizability.
2 Methods
In this section, we describe in detail the novel pipeline we propose for the weakly-supervised diagnosis identification. The pipeline is sketched in Figure 1 and it can be divided into three main steps:
  1. 0
    BPT - Pre-training of a BERT-based model for Italian medical documents.
  2. 1
    SAL - Semi-automatic labelling of the dataset.
  3. 2
    WLC - Classification of discharge letters using the weakly-labelled dataset.
Step 0 (BPT) is described in Section 2.0.1. It aims to build a model able to produce embeddings of Italian medical documents that are exploited by both Step 1 (SAL) and Step 2 (WLC), described in Section 2.0.2 and 2.0.3.
Figure 1: Pipeline schema. Diagram of the full pipeline for weakly-supervised classification of discharge letters
2.0.1 Pre-training of a BERT-based model for Italian medical documents (BPT)
Statistical and machine learning models for textual data require a mathematical representation of the text. There are three main types of possible representations: vocabulary-based, static neural network-based embeddings and contextual neural network-based embeddings. Vocabulary-based representations produce a high dimensional vector whose size is equal to the corpus vocabulary size and whose content is sparse. Their main issue is that they cannot capture the semantic relationships between words, i.e., the fact that different words may have the same or similar meaning, possibly also depending on the context in which they appear. These issues are overcome by neural network-based embeddings. In particular, static word embeddings, such as Word2Vec [32], produce embedding vectors that are permanently fixed for each term. In contrast, contextual embeddings are related to the context in which the word appears, capturing more semantics at the price of requiring an inference through a neural network each time the embedding has to be computed. The state-of-the-art model for contextual embeddings is BERT (Bidirectional Encoder Representations from Transformers) [33], a deep learning model based on the transformer architecture [34]. Different versions of BERT have been developed for different languages and domains. In particular, for the English language, there are various versions of BERT related to the medical domain, such as BioBERT [35] and ClinicalBERT [36].
For the Italian language, the only available versions of BERT are trained on general-domain documents crawled from the web [37]. Thus, we fine-tuned the last Italian version of BERT, Umberto, using the E3C Corpus [38], a multi-lingual dataset composed of documents in the medical domain, particularly clinical narratives from various sources: journal papers, admission tests for specialities in medicine, patient information leaflets for medicines, and abstracts of theses in medical science. The dataset contains documents in five languages; a summary of its composition is provided in Supplementary Material A3. This is the only publicly available dataset that includes Italian medical documents. To utilize the entire dataset, which amounts to 564 MB, we automatically translated the non-Italian documents using Google Translate’s APIs (Step 0.1 in Figure 1). We further pre-trained the Umberto model with this dataset, adopting the usual masked-word prediction task (Step 0.2 in Figure 1). This additional pre-training of the Umberto model over the E3C corpus involved 3 epochs of training (until the validation perplexity did not improve further) with a learning rate of $2\cdot 10^{-5}$ and weight decay of $0.01$, using 20% of the data as a validation set for early stopping.
2.0.2 Semi-automatic labelling pipeline (SAL)
We aim to develop a classification system for discharge letters that can work on an unlabelled dataset. To do so, we aim to construct weak labels that a supervised classification model can subsequently use. This section presents a methodology to construct weak labels by extracting diagnosis strings from discharge letters and clustering them. This allows the creation of groups of diagnoses related to the same disease or symptoms that can be exploited as labels. This step corresponds to Step 1 (SAL) of Figure 1.
Our approach combines regular expressions for extracting diagnosis strings and a customized clustering pipeline based on the BERT model.
In the following paragraphs, we describe in detail these sub-steps.
Extraction of diagnosis strings
The first step, corresponding to block 1.1 in Figure 1, is the extraction of diagnosis strings. Italian discharge letters do not follow a standardized template, as each hospital has its own, and even different departments within the same hospital may personalize the template further. Nevertheless, we have observed that these templates often include a designated section for the diagnosis, where the doctor should report the patient’s admission diagnosis. This diagnosis indication should contain, in most cases, the necessary information to determine the label for the discharge summary. However, there are three main challenges in handling this diagnosis indication:
  1. 1.
    Not all discharge summaries include this indication, as it may be absent or left empty by the physician.
  2. 2.
    Different templates use varying terminologies and position this indication differently in the text.
  3. 3.
    Diagnosis strings are typically written as free text, with the same disease described by different strings and different levels of detail
The first point is the primary reason why our pipeline focuses on classifying the text of a discharge letter rather than just the diagnosis string. To address the second point, we employ regular expressions to detect the portion of text where they are present. Regular expressions provide a powerful and flexible approach for pattern matching, making them suitable for capturing a template field with limited variations, such as the diagnosis indication in our dataset. We reported in Supplementary Material C the expressions we defined to extract the diagnosis strings.
After this step, we have, for the majority of the letters, a string, typically a short sentence, explaining the reasons for the patient’s admission. Due to the third point mentioned before, it is not possible to directly exploit them as labels. To do this, it is necessary to cluster them, as detailed in the following paragraph.
Diagnosis Clustering
Once the diagnosis strings are extracted, we use a tailored clustering pipeline to group them meaningfully, aiming at identifying clusters related to the various types of diseases and symptoms. This step corresponds to block 1.2 in Figure 1 and is expanded in Figure 2.
Figure 2: Clustering section of the pipeline. Detailed diagram of the clustering pipeline for diagnosis strings, corresponding to Step 1.2 of the full pipeline depicted in Figure 1.
Punctuation removal, digits removal and abbreviation expansion are applied to the diagnosis strings (step 1.2.1). This pre-processing aims to facilitate the subsequent construction of a text representation that can be representative of the semantics of the text. In particular, punctuation signs and digits are irrelevant for discriminating between diagnoses and can be safely removed. The problem of abbreviations is a common problem in medical documents: a wide variety of acronyms exist in the medical domain, and their meaning can also be context-dependent. We have extracted all the words with two and three letters from our dataset of diagnosis strings and expanded those that we found to be medical abbreviations.
After this pre-processing, the pipeline needs to construct an embedding of the text that a clustering algorithm will subsequently use. This step (1.2.2) exploits the BERT model we discussed in Section 2.0.1 (BPT).
The BERT-based models produce embedding vectors whose size is 768, making it necessary to apply a dimensionality reduction technique over these embeddings to use them as input for a clustering algorithm (Step 1.2.3).
After this step, a clustering algorithm can be applied to the resulting embedding vectors (Step 1.2.4). We selected HDBSCAN since it provides many relevant advantages: it does not require defining a priori the number of clusters, it does not necessarily cluster all the data, and it can identify clusters of varying shapes and sizes.
To effectively group strings related to different diseases and symptoms, the clustering must prioritize the intra-cluster similarity against the inter-cluster dissimilarity. This has the drawback of creating multiple clusters associated with each disease/symptom that are very compact internally, but that should be merged. A second-level clustering can help merge these clusters that are initially separate, simplifying the identification of the weak labels. To this purpose, a suitable representation for clusters needs to be produced. We employ a keyword representation in which, for each cluster, we extract words that appear with significant frequency in the cluster and do not appear with considerable frequency over the entire dataset (Step 1.2.5). These keywords can be represented with BERT embeddings in the same way as the original diagnosis strings and used for subsequent clustering (Step 1.2.6).
Mapping clusters to diseases
Step 1.3 consists of mapping the clusters to the weak labels that will be used by the classification model. This requires analyzing the diagnosis clusters produced by Step 1.2, using their keyword-based representation (Step 1.2.5). The selection can be fully automated with keyword matching or manual. To automate this step, we propose the following procedure:
  1. 1.
    Definitions: identify a set of definitions for the disease or symptoms of interest. Each definition includes one or more positive keywords (directly related to the disease) and, if necessary, negative keywords (terms indicating alternative diagnoses).
  2. 2.
    Cluster selection: for each cluster, select it if it contains all the positive keywords and none of the negative keywords, according to at least one of the definitions.
This procedure requires minimal manual effort upfront for defining keywords and allows for a more precise selection than other methods, such as using similarity scores between embeddings of definitions and cluster representations.
A keyword-based selection at this step is more straightforward and effective than applying keywords directly to the diagnosis strings extracted in Step 1.1. The original strings are often too heterogeneous, leading to false negatives due to varied terminology and false positives from irrelevant information (e.g., negations, past medical issues). Similarly, manually reviewing the original diagnosis strings is impractical due to their large number.
2.0.3 Classification of discharge letters using the weakly-labelled dataset (WLC)
The classification pipeline that exploits the weakly labelled data produced by Step 1 corresponds to Step 2 in Figure 1 (WLC). A pre-processing step (Step 2.1) is necessary to remove headers and footers that contain information that is non-relevant to our analysis. Supplementary Material B reports details about this. The classification model we adopted is based on the fine-tuned version of the Umberto model that we presented in Section 2.0.1 (BPT). In this case, the transformer-based model is enhanced with a linear classification layer that outputs the probability of having the disease. It is trained with the discharge letters’ text and the weak labels assigned to the letters based on the clusters to which their diagnoses have been assigned (Step 2.2).
Given the limitation of BERT-based models in dealing with texts longer than 512 tokens, when this is the case, we truncated the letter text to the first 512 tokens after pre-processing, possibly including the diagnosis sentences, even if at the end of the letter (in the setting in which we kept it). This is not a relevant limitation, considering that 90% of discharge letters in our dataset are within 512 tokens after pre-processing. In the case of datasets with longer texts, it is possible to chunk them in portions no longer than 512 tokens, process them in parallel and combine their outputs, taking the maximum probability.
2.1 Training and evaluation
The BERT-based classifier is trained with the AdamW optimizer, using a linear learning rate scheduler starting from $10^{-3}$, with a weight decay of 0.01, for six epochs. Initially, only the last classification layer is trained, and after two epochs, the last BERT layer is unfrozen. After four epochs, the last but one BERT layer is unfrozen. These hyperparameters have been selected with grid search.
We cannot perform a precise separate evaluation of all the pipeline steps since we do not have ground truth labels related to diagnosis strings and all their clusters. Because of this, we assess the performance of the whole pipeline on the classification task, implicitly evaluating all the previous steps. For the clustering, we can perform a specific evaluation limited to the clusters related to our bronchiolitis gold labels.
[…]
The classification results are evaluated in terms of precision, recall and F1-score for the positive class and area under the receiver operating characteristic curve (AUC). These metrics are computed for both our weak labels and the gold standard bronchiolitis labels. Evaluating against weak labels allows for an assessment of the classification model’s performance alone, while evaluation against gold standard labels provides a comprehensive view of the pipeline’s effectiveness. We exclude AUC for rule-based classifiers, as they do not produce probability estimates, and we do not assess the fully supervised classifier using weak labels. We compute the metrics with 10-fold stratified cross-validation.
Considering that our weak labels are derived from clusters of diagnosis strings that we extracted from discharge letters, this might bias the classification models into looking only into the diagnosis sections of the letters. To assess the presence of this bias and to verify whether the models can exhibit good performance with the remaining text of the string, we test the models in two settings:
  1. 1.
    With the original letters, possibly including diagnosis strings
  2. 2.
    Removing the diagnosis strings exploited to assign weak labels
Since our dataset contains data from different hospitals and different LHUs, we verify the presence of differences among them. Since the number of cases for each hospital and LHU is limited, we consider those with a minimum number of bronchiolitis cases (15) and measure the results using a Leave One Group Out (LOGO) approach. For each hospital/LHU to be evaluated, we train the model on the data from the other hospitals/LHUs, and we measure the performance on the data from the hospital/LHU of interest.
We also compare results between the groups of pediatric vs non-pediatric ER/Department admissions. In this case, we adopt the same 10-fold stratified cross-validation used to compare different classification models, dividing the test set between pediatric and non-pediatric ERs/Departments at each iteration and computing the metrics on these two groups separately.
3 Results
In this section, we present the results of our pipeline for the bronchiolitis case study, for the different stages that compose it. In particular, Section 3.1 describes the dataset, Section 3.2 presents results related to the extraction of diagnosis strings from discharge letters, Section 3.3, covers the clustering pipeline and Section 3.4 presents the final results on the identification of bronchiolitis.
3.1 Dataset
The dataset comprises 33’176 discharge letters related to children who were admitted to Emergency Rooms and/or hospitalized in the Veneto Region between 2017 and 2020. The dataset covers 44 different hospitals in the region, which can be grouped into 9 Local Health Units (LHUs) and 2 university hospitals. Each discharge letter in the dataset includes the complete text of the letter, along with the date of the letter and anonymized identifiers of the patient. A manual review of discharge letters was conducted to extract relevant diagnoses of bronchiolitis, resulting in 335 positive cases within the dataset. This review was performed by a data manager with a clinical background, with the support of clinicians for uncertain cases. These labels are the gold standard to validate our pipeline, and we do not make direct use of them in the pipeline training.
We identified the records related to pediatric ERs/Departments by looking at the presence of specific keywords (e.g., pediatria, pediatrico, pediatrica) in the header of the letters. This marked 10’340 records (31.17%) as related to pediatric ERs/Departments. Supplementary Material A1 reports additional plots with data distributions among hospitals and LHUs. In particular, ten hospitals cover more than 85% of the records.
The median length of the letters is 4515 characters (1st-3rd quartiles 3240-5806). An example of a fictitious letter is reported in Supplementary Material A2.
3.2 Diagnosis extraction
Our diagnosis extraction module (Step 1.1) successfully extracted a diagnosis sentence from 89% of the discharge letters in the dataset. We extracted 15’038 unique diagnosis sentences, of which 75% occur only once, confirming the need for clustering them in wider groups that can be more easily related to known diseases and symptoms. A few of the extracted strings include the ICD9-CM codes and their standard textual description. Specifically, 737 diagnosis strings (4.9%) contain an ICD9-CM code, covering 3’436 discharge letters (10.4%). The most common strings are the simplest, while more articulate ones have single occurrences. The median length of the extracted diagnoses strings is 49 characters, with 1st and 3rd quartiles of 17 and 156 characters. Among the 335 bronchiolitis-labelled letters, we extracted 138 unique diagnosis strings from 306 letters.
3.3 Diagnosis clustering
In the first clustering level, the 15’038 extracted diagnosis strings were grouped into 590 clusters. Cluster sizes varied significantly, from a large cluster of 270 strings to several smaller clusters containing 5 strings. Table 1 presents the 20 largest clusters, including their sizes and descriptive keywords. The subsequent second-level clustering reduced the number of clusters to 135.
For our case study on bronchiolitis identification, Table 2 details the main clusters related to respiratory infections, including the number of bronchiolitis cases they cover. Four clusters (#1, #2, #3, #5) contain the keyword “bronchiolitis”. Three include a large majority of bronchiolitis cases, whereas the fourth cluster (#5) has only half of its diagnoses labelled as bronchiolitis. These four clusters together cover 65.38% of all bronchiolitis cases. Additionally, 13 clusters include the keyword “bronchospasm”. According to physicians who annotated the data, bronchospasm is associated with bronchiolitis only if accompanied by fever. Only one of these 13 clusters (#4) has both “bronchospasm” and “fever” keywords, and indeed, it contains a high percentage (70%) of bronchiolitis, accounting for an additional 15% of total bronchiolitis cases. Thus, these five clusters are those that, without knowing the gold labels, would be selected to define the weak labels for the training set, according to the two definitions of {bronchiolitis} and {bronchospasm, fever}. Among the respiratory infection clusters, three other clusters contain bronchiolitis cases (#6, #7, #8), but they are in a limited number.
For the second-level clustering, Table 3 presents clusters related to respiratory infections. Some clusters from Table 2 have been merged: two related to mild bronchiolitis, eight related to bronchospasm, and five about generic respiratory infections. The selection of weak labels for training remains unchanged since, looking at their keywords, we select second-level clusters #1, #2, #3, #4 that cover the same records as the clusters we would have chosen at the first level. The main advantage of second-level clustering is the substantial reduction in the number of clusters (approximately 1/5 of the original number of clusters in our dataset), simplifying weak labelling, especially if manual verification is involved.
This cluster selection for weak labels covers 80% of the original bronchiolitis cases. To understand the reason for the false negatives, we reviewed twenty letters labelled as bronchiolitis in the gold standard but not covered by our weak labels. We found 10 cases where the diagnosis pertains to other symptoms or disease (e.g., otitis, dyspnea, fever), but the bronchiolitis is mentioned in the letter, at least as a suspect condition; 4 cases where bronchiolitis is mentioned in the diagnosis alongside other diseases, leading to assignment to a cluster related to the other disease; 3 cases of respiratory issues with no explicit mention of bronchiolitis, respiratory syncytial virus, or bronchospasm and fever; 2 cases with no diagnosis string present; 1 case where a family member has bronchiolitis, but the patient’s diagnosis is of a different disease. These results suggest that the absence of some gold labels in our weak labels primarily arises from the lack of explicit bronchiolitis indications, with only a few cases that could have been detected but were missed. This highlights the importance of using the full discharge letters for a model aiming to identify diseases accurately. Furthermore, there might be a few errors in the gold labels themselves.
We cannot perform a complete evaluation of the clusters produced using our Umberto-E3C model against those produced using the original Umberto model, but we can compare the distribution of gold bronchiolitis labels on the two sets of clusters. Supplementary Material D reports the clusters obtained using the original Umberto model in our clustering pipeline. The number of clusters related to respiratory infections is higher (34 vs 27), and their keywords appear more heterogeneous. Selecting weak labels by applying the same criterion we used on our clusters, we would cover only 47% of bronchiolitis cases with the clusters from the original Umberto model. Moreover, they are less compact, introducing more false negatives in the weak labels.
Table 1: Keyword-based representation of the twenty largest first-level diagnosis clusters
Cluster size Cluster keywords [ITA] Cluster keywords [ENG]
270 flogosi vie phlogosis airways
145 trauma cranico head injury
143 contusione dito mano contusion finger hand
116 gastroenterite verosimile probable gastroenteritis
114 episodio febbre fever episode
94 vie infezione tract infection
93 pronazione dolorosa painful pronation
78 ferita labbro lip wound
75 otite media acuta acute otitis media
70 ferita lacero-contusa lacerocontused wound
66 trauma contusivo blunt trauma
66 broncospasmo flogosi vie bronchospasm pathway inflammation
64 puntura insetto insect bite
62 ferita lacero-contusa lacerocontused wound
59 trauma dx right-handed trauma
58 ustione grado burn degree
54 frattura ferita wound fracture
54 ferita lacero contusa destra right-handed lacerocontused wound
53 insorgenza recente febbre recent onset fever
53 contusione piede destro dito right foot toe contusion
Table 2: Keyword-based representation of the first-level clusters related to respiratory infections, with the % of bronchiolitis cases in each cluster and the % to the total number of cases in the dataset
Cluster # Cluster keywords [ITA] Cluster keywords [ENG]
Within cluster
bronch. %
% of total bronch.
covered
1 bronchiolite lieve mild bronchiolitis 85.19% 6,87%
2 acuta bronchiolite iniziale lieve acute mild initial bronchiolitis 85.14% 44,48%
3 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 74.05% 11,94%
4 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 70,00% 15,02%
5 alimentazione bronchiolite difficolta’ feeding difficulties bronchiolitis 50.00% 2,09%
6 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 12.7% 2,39%
7 broncospasmo infezione respiratoria bronchospasm respiratory infection 6.67% 0,30%
8 broncospasmo disidratazione lieve bronchospasm mild dehydration 1.17% 0,60%
9 flogosi alte vie respiratorie febbre high airways flogosis fever $0$,00%
10 broncospasmo infezione vie respiratorie bronchospasm airways infection $0$,00%
11 acuto broncospasmo acute bronchospasm $0$,00%
12 broncospasmo flogosi vie bronchospasm airways flogosis $0$,00%
13 broncospasmo polmonite bronchospasm pneumonia $0$,00%
14 broncospasmo corso ongoing bronchospasm $0$,00%
15 broncopolmonite sinistra left bronchopneumonia $0$,00%
16 febbricola febbrile infezione respiratoria virosi febrile fever respiratory infection virosis $0$,00%
17 broncospasmo episodio bronchospasm episode $0$,00%
18 infezione alte vie respiratorie high airways infection $0$,00%
19 flogosi infezione alte vie respiratorie flogosis high airways infection $0$,00%
20 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis $0$,00%
21 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm $0$,00%
22 broncospasmo otite bronchospasm otitis $0$,00%
23 broncospasmo parainfettivo bronchospasm parainfective $0$,00%
24 broncospasmo virosi bronchospasm virosis $0$,00%
25 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory $0$,00%
26 crisi epilessia infezione paziente polmonare respiratoria seizure epilepsy respiratory infection lung patient $0$,00%
27 altre infezioni acute vie respiratorie other acute respiratory infection airways $0$,00%
Table 3: Keyword-based representation of the second-level clusters related to respiratory infections, with the % of bronchiolitis cases in each cluster and the % to the total number of cases in the dataset
2nd level cluster # 2nd level clusters keywords [ITA] 2nd level clusters keywords [ENG] 1st level cluster # 1st level clusters keywords [ITA] 1st level clusters keywords [ENG] Within cluster bronch. % % of total bronch. covered 1 bronchiolite lieve mild bronchiolitis 1 acuta bronchiolite iniziale lieve acute mild initial bronchiolitis 85.16 % 51.34 % 2 bronchiolite lieve mild bronchiolitis 2 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 3 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 74.05 % 11.94 % 3 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 4 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 70.00 % 15.02 % 4 bronchiolite difficolta di alimentazione feeding difficulties bronchiolitis 5 bronchiolite difficolta di alimentazione feeding difficulties bronchiolitis 50.00 % 2.09 % 5 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 6 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 12.7 % 2.39 % 6 broncospasmo infezione respiratoria bronchospasm respiratory infection 7 broncospasmo infezione vie respiratorie bronchospasm infection airways 2.9 % 0.30 % 10 broncospasmo infezione respiratoria bronchospasm respiratory infection 7 broncospasmo disidratazione lieve bronchospasm mild dehydration 8 broncospasmo disidratazione lieve bronchospasm mild dehydration 1.17 % 0.60 % 8 broncospasmo bronchospasm 11 acuto broncospasmo acute bronchospasm 0.00 % 0.00 % 12 broncospasmo flogosi vie bronchospasm airways flogosis 13 broncospasmo polmonite bronchospasm pneumonia 14 broncospasmo corso ongoing bronchospasm 22 broncospasmo otite bronchospasm otitis 23 broncospasmo parainfettivo bronchospasm parainfective 24 broncospasmo virosi bronchospasm virosis 17 broncospasmo episodio bronchospasm episode 9 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis 20 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis 0.00 % 0.00 % 10 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm 21 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm 0.00 % 0.00 % 11 infezione vie respiratorie airways infection 18 infezione alte vie respiratorie high airways infection 0.00 % 0.00 % 19 flogosi infezione alte vie respiratorie flogosis high airways infection 0.00 % 0.00 % 27 altre infezioni acute vie respiratorie other acute respiratory infection airways 0.00 % 0.00 % 26 crisi epilessia infezione paziente polmonare respiratoria seizure epilepsy respiratory infection lung patient 0.00 % 0.00 % 16 febbricola febbrile infezione respiratoria virosi febrile fever respiratory infection virosis 0.00 % 0.00 % 12 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory 25 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory 0.00 % 0.00 % 13 alte febbre flogosi respiratorie vie high fever flogosis airways 9 flogosi alte vie respiratorie febbre high airways flogosis fever 0.00 % 0.00 % 14 broncopolmonite sinistra left bronchopneumonia 15 broncopolmonite sinistra left bronchopneumonia 0.00 % 0.00 %
3.4 Disease classification
[…]
When comparing the classification results on weak labels (W) and gold labels (G), we observe a reduction in precision, affecting both the F1-Score and the AUC. However, recall is higher on gold labels, at least for BERT-based models, whereas simpler models also showed a decrease in recall.
The BERT-based models show minimal differences between settings with and without diagnosis string, indicating that the information needed to identify the diagnosis of bronchiolitis is present in the entire letter text, and these models can effectively recognize it.
As expected, the fully supervised model demonstrated better overall performance, though differences are modest, particularly in recall, which is lower on average than our best weakly-supervised model.
In Table 5, we assess the differences in the results of our best model when selecting different subsets of clusters for the weak labels. In particular, we exclude one of the 1st level clusters we selected at a time. Excluding clusters #5 or #1 does not affect the results, which is reasonable considering their smaller size. Excluding cluster #2 determines a drop in the performance and an increase in the variance, which is expected since it accounts for almost half of the bronchiolitis cases. Removing clusters #3 or #4 has a negative effect, but limited. These results confirm the solidity of our weak labels and the robustness of the model with respect to small changes in them.
Table 6 compares metrics at the hospital level, revealing notable differences. Some hospitals, such as H-0 and H-3, are more represented in the dataset, which might explain their superior performance. However, hospitals with a similar number of records show considerable performance variability, suggesting that reporting practices in discharge letters might influence disease detection. A comparison between records from pediatric vs non-pediatric ER/Departments revealed significant differences (gold AUC $80.89\pm 6.37$ vs $71.07\pm 4.57$, see Supplementary Material E for details), but there are also hospitals with pediatric ER and low performances, such as H-4. Similar findings are reported in Supplementary Material E for the different LHUs.
Table 4: 10-fold CV results for classification, comparing with rule-based baselines, fully supervised approaches and other classification modules in our pipeline. Std. Dev. in parenthesis. G = gold labels, W = weak labels.
With diagnosis string
Model
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
Umberto-E3C 82.30 (8.94) 75.46 (6.25) 77.33 (4.06) 82.30 (4.99) 72.75 (7.50) 84.59 (6.48) 75.06 (3.72) 77.68 (5.73)
Umberto 80.67 (5.99) 73.48 (7.12) 76.91 (5.97) 82.56 (4.79) 62.29 (6.24) 75.99 (4.69) 68.46 (2.77) 69.82 (4.54)
LSTM 68.57 (6.48) 64.97 (11.96) 65.54 (5.66) 67.50 (3.99) 57.39 (11.37) 60.86 (13.67) 55.08 (8.07) 63.89 (5.28)
Rule-based full text 70.18 87.05 77.70 - 65.57 68.76 67.13 -
Rule-based diagnosis 73.88 86.38 79.64 - 74.61 67.42 70.83 -
Umberto-E3C Supervised - - - - 80.30 (7.23) 81.52 (6.70) 80.91 (2.81) 86.45 (2.82)
Without diagnosis string
Model
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
Umberto-E3C 84.34 (5.19) 73.04 (10.53) 75.15 (4.27) 84.91 (3.50) 70.15 (6.40) 79.00 (5.82) 74.28 (4.10) 77.56 (3.11)
Umberto 80.01 (6.84) 68.68 (10.55) 73.90 (5.36) 80.81 (6.24) 61.41 (6.09) 76.69 (4.22) 68.18 (3.62) 70.49 (3.89)
LSTM 64.98 (6.89) 56.21 (18.88) 58.80 (13.52) 63.00 (9.24) 50.14 (8.31) 63.97 (13.52) 53.03 (9.69) 59.02 (7.36)
Rule-based full text 65.54 79.16 71.69 - 62.33 60.05 61.10 -
Umberto-E3C Supervised - - - - 75.56 (5.08) 82.55 (9.14) 78.90 (4.54) 84.05 (4.26)
Table 5: 10-fold CV results for classification, considering different groups of 1st level clusters for weak labels. Std. Dev. in parenthesis. G = gold labels, W = weak labels
Clusters for weak labels
(1st level)
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
#1. #2. #3. #4 78.39 (7.04) 73.35 (6.21) 75.79 (4.65) 83.02 (4.16) 70.75 (6.48) 81.27 (6.99) 75.65 (4.29) 75.49 (5.57)
#1. #2. #4. #5 73.69 (8.89) 83.20 (9.16) 78.16 (7.80) 80.71 (7.01) 68.42 (8.70) 75.09 (8.58) 71.60 (7.26) 72.67 (6.87)
#2. #3. #4. #5 83.49 (7.87) 74.66 (6.88) 78.82 (4.26) 84.02 (4.43) 68.54 (6.30) 82.35 (4.94) 74.81 (3.82) 75.83 (5.22)
#1. #3. #4. #5 93.73 (6.85) 78.11 (8.47) 85.19 (7.06) 90.91 (7.35) 57.29 (12.97) 59.03 (15.94) 58.15 (12.45) 68.90 (10.04)
#1. #2. #3. #5 80.45 (8.90) 74.62 (8.57) 77.43 (5.76) 85.95 (5.29) 67.86 (6.36) 73.42 (8.64) 70.53 (5.33) 70.22 (3.10)
Table 6: 10-fold CV results for classification, comparing different hospitals (only hospitals with more than 15 bronchiolitis cases). G = gold labels, W = weak labels.
Hospital Dataset %
Pediatric ER
Bronchiolitis
P [%]
R [%]
F1 [%]
AUC [%]
P [%]
R [%]
F1 [%]
AUC [%]
H-0 31.41 87.30 0.76 77.46 78.57 78.01 90.40 81.40 63.64 71.43 86.56
H-1 14.56 0.00 1.08 63.64 71.79 79.43 85.32 66.67 66.67 66.67 77.95
H-2 11.63 1.00 0.78 80.00 72.73 76.19 92.16 50.00 71.43 58.82 83.98
H-3 10.23 12.00 1.71 88.68 88.68 88.68 89.45 81.40 81.40 81.40 88.12
H-4 7.43 22.00 1.22 92.31 35.29 51.06 64.32 84.21 37.21 51.61 57.34
H-6 2.59 9.40 2.09 100.00 31.25 47.62 65.63 100.00 33.33 50.00 70.48
4 Discussion
This study introduces a novel pipeline that shows how weakly-supervised learning can be effectively conducted in the clinical natural language processing domain. This is the first study to address the problem of diagnosis identification from discharge letters in the Italian language, and while we validated the method using a case study on bronchiolitis, the pipeline is designed to be applicable to any disease. Our case study shows excellent results, outperforming existing unsupervised approaches, and underscores the potential for broader application in clinical settings.
We observed minimal performance differences between the metrics computed on weak and gold labels (1-2 % of F1-Score, 5-7 % of AUC), confirming that the first step of our pipeline (SAL) effectively identifies relevant cases. Despite the inevitable noise that the weak labels introduce, this is limited enough to allow the model to perform well when evaluated against gold labels. Additionally, the variations between settings where we retained the diagnosis strings and those where we removed them are negligible. This indicates that the model is effectively learning from the entire text of the discharge letters, making it particularly advantageous in situations where automatic extraction of diagnosis strings is not possible. Our manual exploration of some weak labels further revealed that diagnosis strings do not always mention the disease responsible for hospitalization. In some cases, they only refer to symptoms, and in cases of multiple diseases, not all may be reported. These findings highlight the necessity for a classification model based on the complete text of discharge letters, as the one we developed.
In contrast, rule-based approaches exhibited poor performance, mainly due to the fact that diseases like bronchiolitis are not always explicitly mentioned by name in the text. Even when mentioned, they might refer to discarded hypotheses or contacts with affected individuals, leading to false positives. While a more sophisticated rule-based system could address some of these issues, such approaches lack the scalability and generalizability of our pipeline. Conversely, our pipeline can be easily adapted to other diseases, which makes it a versatile tool for clinical NLP.
Our fine-tuned version of the Umberto model showed the best results both in the classification and the clustering modules, demonstrating a better understanding of the Italian medical domain when compared with the general-domain version.
Compared to a fully supervised method, our weakly-supervised pipeline maintained competitive performance, with only a limited and acceptable loss ($\sim 5\%$ of F1-Score, $\sim 6\%$ of AUC). This is particularly notable given the significant time and effort required for manual annotation in fully supervised setups. The pipeline has also shown robustness to the cluster selection for the weak labels. These results underscore the potential of weakly-supervised approaches in clinical natural language processing, a domain where the scarcity of labelled data is the most relevant challenge.
While our results are promising, we must acknowledge some limitations in our study that could be addressed in future work. The absence of a gold standard in labelling is a common challenge in clinical NLP, which underscores the value of weakly-supervised methods like ours. Our gold labels for bronchiolitis were manually assigned by individual physicians, making it impossible to verify inter-annotator agreement and increasing the risk of errors.
The multi-centric nature of our dataset, though a strength in its diversity, also presents an imbalance in the number of records from different hospitals and Local Health Units. Our analysis highlighted relevant performance differences among hospitals and LHUs, warranting further investigation. Testing the pipeline on a more balanced dataset would strengthen conclusions about these performance differences. Moreover, this could facilitate the development of a hierarchical transformer-based classification model that considers hospital/LHU context in the input, potentially improving model performance.
We also observed relevant differences between pediatric vs. non-pediatric ERs/Departments. Regarding this, we need to consider that pediatric admissions were identified using specific keywords. This method could miss some pediatric cases, although such instances should be limited. Future work could use data that directly includes this information as a structured field, ensuring its accuracy and completeness.
Although our case study focused on bronchiolitis, the diagnostic clusters we identified appear capable of recognizing a broad range of diseases. Validating the pipeline on other diseases would further demonstrate its robustness and generalizability. Developing a multi-disease model, particularly for conditions that share similar symptoms, could provide broader applicability and enhance the model’s utility. Tailored rule-based models often fail in such scenarios, making a machine learning-based approach more effective.
Another important direction for future research is enhancing the interpretability of our model. Identifying the portions of the discharge letter that are most relevant to the model’s output would help clinicians understand and trust the model’s decisions. Techniques related to the area of explainable AI could be employed to enhance interpretability, ensuring the model’s reasoning process is transparent and understandable. This would also help mitigate biases that could potentially be present in the training data, further enhancing the model’s reliability.
5 Conclusions
We have developed and validated the first weakly-supervised pipeline to analyze discharge letters in the Italian language, with a case study focusing on bronchiolitis. The pipeline achieved excellent results, outperforming rule-based approaches. Our fine-tuned BERT-based model showed superior performance in both the clustering and the classification module compared to a general-domain BERT model. We also assessed significant differences between hospitals/LHUs and pediatric vs. non-pediatric ERs, which suggest further investigations.
However, while the case study showcased the method’s potential, the pipeline itself is designed to be adaptable to any disease, making it a valuable tool for researchers interested in extracting diagnoses from unstructured data and for epidemiological monitoring of diseases.
Despite a small performance gap compared to fully supervised approaches, the weakly-supervised pipeline addresses real-world constraints, particularly the scarcity of labelled data in clinical NLP. Transformer-based models showed strong performance even without gold-labelled data, underscoring the viability of weakly-supervised methods in overcoming data limitations in the medical domain.
Declaration of competing interests
The authors declare that they have no competing interests.
Data statement
Data used in this study are collected and accessed through Pedianet, an Italian network of paediatricians. Data anonymization is performed in accordance with Italian regulations. Each anonymized record is assigned a unique numerical identifier. The data is then transmitted monthly to a centralized database located in Padua, Italy, for validation purposes. Inclusion in the Pedianet database is voluntary and parents/legal guardians must provide consent for their children’s anonymized data to be used for research purposes. Ethical approval of the study and access to the database were granted by the Internal Scientific Committee of So.Se.Te. Srl, the legal owner of Pedianet.
Data used in this study are not publicly available due to their confidential nature. The code developed for their analysis is available by the authors upon request.
CRediT authorship contribution statement
Vittorio Torri: Methodology, Software, Investigation, Writing - Original Draft Elisa Barbieri: Conceptualization, Data Curation, Writing - Review & Editing Anna Cantarutti: Conceptualization, Data Curation, Writing - Review & Editing Carlo Giaquinto: Conceptualization, Supervision Francesca Ieva: Conceptualization, Methodology, Writing - Review & Editing, Supervision
Acknowledgments
The present research has been partially supported by MUR, grant Dipartimento di Eccellenza 2023-2027.
Elisa Barbieri declares that her position is funded within the project “INF-ACT - One Health Basic and translational Research Actions addressing Unmet Needs on Emerging Infectious Diseases”, ID MUR PE_00000007, under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2 – CUP C93C22005170007.
The authors thank all the family paediatricians collaborating in Pedianet: Eva Alfieri, Michela Alfiero Bordigato, Angelo Alongi, Biagio Amoroso, Rosaria Ancarola, Barbara Andreola, Giampaolo Anese, Roberta Angelini, Maria Grazia Apostolo, Giovanna Argo, Giovanni Avarello, Lucia Azzoni, Maria Carolina Barbazza, Patrizia Barbieri, Gabriele Belluzzi, Eleonora Benetti, Filippo Biasci, Franca Boe, Stefano Bollettini, Francesco Bonaiuto, Anna Maria Bontempelli, Matteo Bonza, Sara Bozzetto, Andrea Bruna, Ivana Brusaterra, Massimo Caccini, Laura Calì, Sonia Camposilvan, Laura Cantalupi, Luigi Cantarutti, Chiara Cardarelli, Giovanna Carli, Sylvia Carnazza, Massimo Castaldo, Stefano Castelli, Monica Cavedagni, Giuseppe Egidio Cera, Chiara Chillemi, Francesca Cichello, Giuseppe Cicione, Carla Ciscato, Mariangela Clerici Schoeller, Samuele Cocchiola, Giuseppe Collacciani, Valeria Conte, Roberta Corro’, Rosaria Costagliola, Nicola Costanzo, Sandra Cozzani, Giancarlo Cuboni, Giorgia Curia, Caterina D’alia, Vito Francesco D’Amanti, Antonio D’Avino, Roberto De Clara, Lorenzo De Giovanni, Annamaria De Marchi, Gigliola Del Ponte, Tiziana Di Giampietro, Giuseppe Di Mauro, Giuseppe Di Santo, Piero Di Saverio, Mattea Dieli, Marco Dolci, Mattia Doria, Dania El Mazloum, Maria Carmen Fadda, Pietro Falco, Mario Fama, Marco Faraci, Maria Immacolata Farina, Tania Favilli, Mariagrazia Federico, Michele Felice, Maurizio Ferraiuolo, Michele Ferretti, Mauro Gabriele Ferretti, Paolo Forcina, Patrizia Foti, Luisa Freo, Ezio Frison, Fabrizio Fusco, Giovanni Gallo, Roberto Gallo, Andrea Galvagno, Alberta Gentili, Pierfrancesco Gentilucci, Giuliana Giampaolo, Francesco Gianfredi, Isabella Giuseppin, Laura Gnesi, Costantino Gobbi, Renza Granzon, Mauro Grelloni, Mirco Grugnetti, Antonina Isca, Urania Elisabetta Lagrasta, Maria Rosaria Letta, Giuseppe Lietti, Cinzia Lista, Ricciardo Lucantonio, Francesco Luise, Enrico Marano, Francesca Marine, Lorenzo Mariniello, Gabriella Marostica, Sergio Masotti, Stefano Meneghetti, Massimo Milani, Stella Vittoria Milone, Donatella Moggia, Angela Maria Monteleone, Pierangela Mussinu, Anna Naccari, Immacolata Naso, Flavia Nicoloso, Cristina Novarini, Laura Maria Olimpi, Riccardo Ongaro, Maria Maddalena Palma, Angela Pasinato, Andrea Passarella, Pasquale Pazzola, Monica Perin, Danilo Perri, Silvana Rosa Pescosolido, Giovanni Petrazzuoli, Giuseppe Petrotto, Patrizia Picco, Ambrogina Pirola, Lorena Pisanello, Daniele Pittarello, Elena Porro, Antonino Puma, Maria Paola Puocci, Andrea Righetti, Rosaria Rizzari, Cristiano Rosafio, Paolo Rosas, Bruno Ruffato, Lucia Ruggieri, Annamaria Ruscitti, Annarita Russo, Pietro Salamone, Daniela Sambugaro, Luigi Saretta, Vittoria Sarno, Valentina Savio, Nico Maria Sciolla, Rossella Semenzato, Paolo Senesi, Carla Silvan, Giorgia Soldà, Valter Spanevello, Sabrina Spedale, Francesco Speranza, Sara Stefani, Francesco Storelli, Paolo Tambaro, Giacomo Toffol, Gabriele Tonelli, Silvia Tulone, Angelo Giuseppe Tummarello, Cristina Vallongo, Sergio Venditti, Maria Grazia Vitale, Concetta Volpe, Francescopaolo Volpe, Aldo Vozzi, Giulia Zanon, Maria Luisa Zuccolo.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT to improve the readability and clarity of some sentences. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.

### v2

@@
-Weakly-supervised diagnosis identification from Italian discharge letters
+Automatic identification of diagnosis from hospital discharge letters via weakly-supervised Natural Language Processing
 Abstract
-Objective: Identifying the diagnosis that determines a patient’s hospitalization or Emergency Room admission is crucial for cohort selection and epidemiological analyses.
-Discharge letters are the only type of data consistently produced across hospitals.
-Recognizing diseases from discharge letters is a classic document classification problem, typically requiring supervised learning.
-However, manual annotation of large datasets of discharge letters is uncommon since it is extremely time-consuming.
-We propose a novel weakly-supervised pipeline to recognize diseases from Italian discharge letters.
-Methods: Our Natural Language Processing pipeline is based on a fine-tuned version of the Italian Umberto model.
-The pipeline extracts diagnosis-related sentences from a subset of letters and applies a two-level clustering using the embeddings generated by the fine-tuned Umberto model.
-These clusters are summarized and those mapped to the diseases of interest are selected as weak labels.
-Finally, the same BERT-based model is trained using these weak labels to detect the targeted diseases.
-Results: Results on a case study related to the identification of bronchiolitis in a dataset of 33’176 Italian discharge letters from multiple hospitals in the Veneto Region demonstrate the potential of our method, with an AUC of 77.7 % and an F1-Score of 75.1 % on manually annotated labels, improving compared to other non-supervised methods and with a limited loss compared to fully supervised methods.
-Results are robust to the cluster selection and the identified clusters highlight the potential to recognize a variety of diseases.
-Conclusions: This study demonstrates the feasibility of successfully addressing the challenge of diagnosis identification from Italian discharge letters in the absence of labelled data.
-Our weakly-supervised pipeline showed strong performance and robustness, and its flexibility allows for easy adaptation to various diseases.
-This approach offers a scalable solution for clinical text classification, significantly reducing the need for manual annotation while maintaining good accuracy.
+Identifying patient diagnoses from hospital discharge letters is essential to enable large-scale cohort selection and epidemiological research, but traditional supervised learning approaches rely on extensive manual annotation, which is often impractical for large textual datasets.
+In this study, we present a novel weakly supervised Natural Language Processing pipeline designed to classify Italian discharge letters without requiring manual labelling.
+After extracting diagnosis-related sentences, the method leverages a transformer-based model with an additional pre-training on Italian medical documents to generate semantic embeddings.
+A two-level clustering procedure is then applied to these embeddings, and the resulting clusters are mapped to the diseases of interest to derive weak labels for a subset of data, which are subsequently used to train a transformer-based classifier.
+We evaluate the approach on a real-world case study focused on detecting bronchiolitis in a corpus of 33,176 Italian discharge letters of children admitted to 44 emergency rooms or hospitals in the Veneto Region between 2017 and 2020.
+The pipeline achieves an area under the curve (AUC) of 77.68% ($\pm 4.30\%)$ and an F1-score of 78.14% ($\pm 4.89\%$) against manually annotated data.
+Its performance surpasses other unsupervised methods and approaches that of fully supervised models, while maintaining robustness to cluster selection and a promising generalizability across different disease types.
+It allows saving approximately 3 minutes of expert time per discharge letter, corresponding to more than 1,500 hours for a dataset of the size used in our case study.
+This study demonstrates the feasibility of a weakly-supervised strategy for identifying diagnoses from Italian discharge letters.
+The pipeline achieves strong performance, is adaptable to various diseases, and offers a scalable solution for clinical text classification, reducing the need for manual annotation while maintaining reliable accuracy.
 keywords:
@@
 Electronic Health Records contain a vast and continuously growing amount of patient information, but a significant portion of this data remains unstructured in formats like clinical notes, reports, discharge letters and referrals.
-Despite containing valuable information, this unstructured data is often underutilized in the downstream analysis due to their complexity [1, 2].
-While Natural Language Processing (NLP) has made significant advancements in extracting and classifying information from text [3, 4], applying these techniques to healthcare data presents additional challenges, such as a specialized lexicon, a high prevalence of ambiguous abbreviations, frequent misspellings, and a lack of standard templates across hospitals [5, 6].
-Additionally, most NLP tools are developed for high-resource languages, making them less effective for languages like Italian, where linguistic resources are scarce [7].
-A critical piece of information in medical documents is the reason for a patient’s hospital admission or discharge, typically a disease diagnosis or a description of the patient’s symptoms.
-This information is critical for both cohort selection and epidemiological analysis.
-Although an international code system for diseases and symptoms, the ICD-CM [8], exists, physicians often prefer to provide a textual description of their diagnoses.
-Textual descriptions are faster to write, more detailed, and are often used because many medical document templates do not include structured fields for disease codes [9, 10].
-Even when present, it has been observed that ICD codes are not completely reliable [11, 12, 13].
-Consequently, identifying patients admitted for specific diseases or symptoms often requires manual inspection of these texts, which is labour-intensive and impractical at scale.
-In Italy, every hospitalization or Emergency Room (ER) admission generates a discharge letter.
-Depending on the type of examinations and the hospital, more structured data may also be recorded.
-A discharge letter should always contain information about the diseases that affect the patient, but the complexity of these documents and the lack of standard templates pose a significant challenge for automatic data extraction.
-To address this issue, this study aims to develop an NLP pipeline capable of extracting structured information from the unstructured text of Italian discharge letters.
-Traditional approaches to this problem typically frame it as a supervised classification task, which requires a substantial amount of annotated data to perform well [14].
-However, annotating thousands of discharge letters is time-consuming, especially for clinicians who are already overworked.
-This highlights a significant gap in the field: the need for methods that can effectively utilize large, unannotated datasets without the burden of manual labelling.
-While there have been a few studies on NLP applied to Italian healthcare documents, none have focused specifically on discharge letters.
+Despite containing valuable information, this unstructured data is often underutilized in the downstream analysis due to their complexity and heterogeneity [tayefi2021challenges, cowie2017electronic].
+While Natural Language Processing (NLP) has made significant advancements in extracting and classifying information from text [nadkarni2011natural, khurana2023natural], applying these techniques to clinical narratives remain challenging.
+Medical documents are characterized by a highly specialized lexicon, with ambiguous abbreviations and frequent misspellings, a lack of standard templates across hospitals and a limited amount of manually labelled data [hossain2023natural, friedman2014natural, iroju2015systematic].
+Additionally, most NLP tools are optimized for high-resource languages such as English, limiting their transferability to languages like Italian, where datasets and domain-specific linguistic resources are scarce [neveol2018clinical].
+Among the various types of clinical documents, hospital discharge letters play a particularly important role: they summarize the patient’s condition, diagnoses, and treatments, representing a relevant source of diagnostic information for both clinical follow-up and population-based studies.
+Although international standard such as the ICD-CM [world2016international] exists for encoding diagnoses, physicians often prefer to provide textual descriptions of their diagnoses because they are faster to write, more expressive, and often the only option available in document templates [walsh2004clinician, millares2021consultation].
+Even when ICD-CM codes are included, previous research has shown that they can be incomplete or inaccurate [walsh2018icd, cozzolino2018accuracy, cipparone2015inaccuracy].
+Automatic methods capable of identifying diagnoses directly from unstructured discharge letters are therefore crucial for cohort selection and epidemiological research [dong2022automated], reducing the need for manual review, which is labor-intensive and impractical at scale.
+In Italy, every hospital or emergency room (ER) admission generates a discharge letter that should contain at least one diagnosis describing the reason for admission or discharge.
+However, these letters vary widely in format and terminology, often mixing narrative and structured components.
+Developing robust NLP systems for such data is challenging: traditional approaches rely on supervised classification models that require a substantial amount of manually annotated data [mehrafarin2022importance], an effort that is costly and time-consuming for clinicians.
+This highlights a significant gap in the field: the need for methods that can effectively exploit large, unannotated datasets to extract clinically meaningful information without the burden of manual labelling.
+Previous studies applying NLP to Italian clinical texts have not addressed discharge letters.
 Lanera et al.
-[15, 16] worked on supervised models to identify cases of varicella, using textual data from specific children’s health project dataset.
-Similarly, Sciannameo et al.
-[17] used a BERT-based model to predict rehospitalizations for elderly people from the text of medication prescriptions and hospitalization diagnoses.
+[lanera2020use, lanera2022deep] worked on supervised models to identify varicella cases from notes of family pediatricians, while Sciannameo et al.
+[veronica2022deep] used a BERT-based model to predict rehospitalizations for elderly patients from medication prescriptions and hospitalization records.
 Hammami et al.
-[18] classified pathology reports for cancer morphology, developing a rule-based system.
-Other studies have concentrated on extracting events and entities, such as Viani et al.
-[19, 20], who analyzed cardiology reports using ontologies and recurrent neural networks.
-Significant research on discharge letters has been done in the English language, often using the MIMIC dataset [21, 22] (see [23, 24, 25, 26] for some examples).
-Research has also been conducted on discharge letters in other under-resourced languages, such as Bulgarian [27, 28], Dutch [29] and Swedish [30].
-All these studies have relied either on large annotated datasets or on the development of complex, problem-specific rule-based systems.
-In contrast, our approach eliminates the dependency on manually annotated data and offers a more generalizable solution than traditional rule-based systems.
-The main contribution of this paper is the first NLP-based pipeline for the weakly-supervised classification of Italian discharge letters, offering a scalable solution to a long-standing challenge in healthcare data management.
-The results presented in this paper pertain to a clinically relevant case study aimed at identifying bronchiolitis, one of the most relevant diseases that can determine hospitalizations in young children [31].
-An automated model for this purpose can significantly aid in patient cohort selection for research, reducing the manual workload for physicians.
-The dataset for this case study comprises 33’176 discharge letters related to children from 44 different hospitals in the Veneto region, one of the most populated Italian regions.
-The rest of the paper is structured as follows: in Section 2, we discuss the methodology we developed and the data of our case study; in Section 3, we present the results; in Section 4 we discuss them, together with possible future developments, and in Section 5 we summarize the main findings.
-Statement of significance
-Problem Identifying diagnoses from discharge letters is relevant for both cohort selection and epidemiological analysis but requires large annotated datasets, rarely available in this domain.
-What is Already Known Supervised learning methods for diagnosis identification are well-studied, with state-of-the-art methods based on transformers.
-Unsupervised methods are mainly rule-based algorithms that lack generalizability.
-What this Paper Adds This paper introduces a weakly-supervised NLP pipeline that eliminates the need for manual annotations, offering a scalable solution for diagnosis identification.
-The approach is validated on a case study with a large dataset of Italian discharge letters and demonstrates strong performance and potential generalizability.
+[hammami2021automated] classified pathology reports for cancer morphology using a rule-based approach, and Viani et al.
+[viani2018information, viani2019supervised] extracted entities from cardiology reports through ontologies and recurrent neural networks.
+Most international work on discharge letters has been conducted in English using resources such as the MIMIC database [johnson2016mimic, johnson2023mimic] (see [edin2023automated, li2020icd, li2018automated, falter2024using] for some examples), but research has also been conducted in other European languages with limited resources, such as Bulgarian [boytcheva2011automatic, velichkov2020automatic], Dutch [bagheri2020automatic, van2024using], Czech [lenc2024czech] and Swedish [remmer2021multi].
+These studies, however, all relied on large annotated datasets, making them difficult to reproduce in settings where manual labelling is infeasible.
+Recent work has explored weakly-supervised approaches to mitigate the need for large annotated clinical corpora.
+Most of these methods rely on task-specific heuristics, such as disease-related keywords, regular expressions, or the transfer of labels from structured data like diagnostic codes or medication records [sanyal2022weakly, wang2019clinical, cusick2021using, wang2022leveraging, humbert2022strategies].
+While effective within narrowly defined clinical settings, such strategies are difficult to extend to new diseases, domains, or institutions, as they depend on predefined lexical patterns or on the availability of reliable structured fields.
+To overcome these limitations, we propose a weakly-supervised NLP pipeline for automatically identifying patient diagnoses from Italian hospital discharge letters.
+Our approach eliminates the dependency on manually annotated data and offers a more generalizable solution than previous rule-based methods.
+The pipeline first identifies diagnosis-related sentences from a subset of documents and represents them through contextual embeddings generated by a transformer model, previously pre-trained with public data in Italian related to the medical domain.
+These embeddings are then grouped through a two-level clustering procedure, after which the resulting clusters are semantically summarized and mapped to diseases of interest to produce weak labels.
+The weakly labelled data are finally used to train a transformer-based classifier capable of recognizing those diseases directly from discharge letters.
+This design enables large-scale, disease-agnostic classification of clinical narratives while avoiding the costs of manual labelling.
+We evaluate the proposed framework through a real-world case study focused on bronchiolitis, one of the leading causes of hospitalizations in young children [zorc2010bronchiolitis].
+The dataset includes 33,176 discharge letters of pediatric patients admitted to 44 emergency rooms or hospitals across the Veneto Region (Italy) between 2017 and 2020, collected through the Pedianet network [cantarutti2021pedianet].
+To the best of our knowledge, there are no prior studies addressing the identification of bronchiolitis directly from medical text, whereas existing works have relied on structured data for diagnosis prediction [durani2008clinical, gebremedhin2022developing, vartiainen2023risk].
+In our evaluation, we test the weakly-supervised pipeline against manually labelled data, considering different classification backbones and comparing its performance with both fully unsupervised and fully supervised approaches.
+The proposed method achieves competitive performance, outperforming other unsupervised techniques and showing only a limited gap compared with supervised classifiers.
+Its performance is robust to weak labels selection, and the analysis of the intermediate steps indicates a strong potential for generalization across diseases.
 2 Methods
-In this section, we describe in detail the novel pipeline we propose for the weakly-supervised diagnosis identification.
-The pipeline is sketched in Figure 1 and it can be divided into three main steps:
+Figure 1: a.
+Diagram of the full pipeline for weakly-supervised classification of discharge letters.
+Step 0 (TPT) consists in additional pre-training of a transformer-based model on a dataset of medical documents in Italian, including documents auomatically translated into Italian.
+Step 1 (SAL) is the core of the pipeline, consisting in the automatic extraction of diagnosis strings from the letters, their clustering (exploiting the model from TPT) and the mapping of the clusters to the disease labels.
+This procedure leads to a weakly-labelled subset of data that is used in Step 2 (WLC) to train a weakly-supervised classificaton model.
+b.
+Detailed diagram of the clustering step for diagnosis strings, corresponding to Step 1.2 of the full pipeline.
+The extracted diagnoses are embedded with a transformer-based model, and, after a dimensionality reduction step, clustered.
+Clusters are summarized with relevant keywords and the clustering is repeated on these clusters’ representations.
+In this section, we describe in detail the pipeline we propose for the weakly-supervised diagnosis identification. The pipeline, depicted in Figure 1, can be divided into three main steps:
   1. 0
-    BPT - Pre-training of a BERT-based model for Italian medical documents.
+    TPT - Pre-Training of a Transformer-based model for Italian medical documents.
   2. 1
@@
     WLC - Classification of discharge letters using the weakly-labelled dataset.
-Step 0 (BPT) is described in Section 2.0.1.
-It aims to build a model able to produce embeddings of Italian medical documents that are exploited by both Step 1 (SAL) and Step 2 (WLC), described in Section 2.0.2 and 2.0.3.
-Figure 1: Pipeline schema. Diagram of the full pipeline for weakly-supervised classification of discharge letters
-2.0.1 Pre-training of a BERT-based model for Italian medical documents (BPT)
+The next sections describe in details the three steps.
+2.1 Pre-training of a Transformer-based model for Italian medical documents (TPT)
 Statistical and machine learning models for textual data require a mathematical representation of the text.
@@
 These issues are overcome by neural network-based embeddings.
-In particular, static word embeddings, such as Word2Vec [32], produce embedding vectors that are permanently fixed for each term.
+In particular, static word embeddings, such as Word2Vec [mikolov2013efficient], produce embedding vectors that are permanently fixed for each term.
 In contrast, contextual embeddings are related to the context in which the word appears, capturing more semantics at the price of requiring an inference through a neural network each time the embedding has to be computed.
-The state-of-the-art model for contextual embeddings is BERT (Bidirectional Encoder Representations from Transformers) [33], a deep learning model based on the transformer architecture [34].
-Different versions of BERT have been developed for different languages and domains.
-In particular, for the English language, there are various versions of BERT related to the medical domain, such as BioBERT [35] and ClinicalBERT [36].
-For the Italian language, the only available versions of BERT are trained on general-domain documents crawled from the web [37].
-Thus, we fine-tuned the last Italian version of BERT, Umberto, using the E3C Corpus [38], a multi-lingual dataset composed of documents in the medical domain, particularly clinical narratives from various sources: journal papers, admission tests for specialities in medicine, patient information leaflets for medicines, and abstracts of theses in medical science.
-The dataset contains documents in five languages; a summary of its composition is provided in Supplementary Material A3.
+The state-of-the-art models for generating contextual embeddings are based on the transformer architecture [vaswani2017attention].
+Among them, BERT (Bidirectional Encoder Representations from Transformers) [devlin2019bert] and its subsequent variants, such as RoBERTa [liu2019roberta], have become foundational for a wide range of NLP tasks.
+Different versions of these models have been further specialized for specific languages and domains.
+In particular, for the English language, there are multiple models for the biomedical domain, such as BioBERT [lee2020biobert] and ClinicalBERT [alsentzer2019publicly].
+For the Italian language, available transformer models equivalent to BERT are mostly trained on general-domain web data [tamburini2020bertology], with limited exposure to biomedical or clinical terminology.
+The E3C corpus [magnini2020e3c] is a multilingual resource that includes Italian documents from diverse medical sources, such as journal articles, specialty-school admission tests, patient information leaflets, and medical-science thesis abstracts, providing valuable clinical context for domain adaptation.
+The dataset contains documents in five languages; a summary of its composition is provided in A.3.
 This is the only publicly available dataset that includes Italian medical documents.
+With the aim of better capturing domain-specific semantics, we used this data to perform an additional pre-training of the transformer-based models we included in our pipeline (Step 0.2 in Figure 1).
 To utilize the entire dataset, which amounts to 564 MB, we automatically translated the non-Italian documents using Google Translate’s APIs (Step 0.1 in Figure 1).
-We further pre-trained the Umberto model with this dataset, adopting the usual masked-word prediction task (Step 0.2 in Figure 1).
-This additional pre-training of the Umberto model over the E3C corpus involved 3 epochs of training (until the validation perplexity did not improve further) with a learning rate of $2\cdot 10^{-5}$ and weight decay of $0.01$, using 20% of the data as a validation set for early stopping.
-2.0.2 Semi-automatic labelling pipeline (SAL)
-We aim to develop a classification system for discharge letters that can work on an unlabelled dataset.
-To do so, we aim to construct weak labels that a supervised classification model can subsequently use.
-This section presents a methodology to construct weak labels by extracting diagnosis strings from discharge letters and clustering them.
+2.2 Semi-automatic labelling pipeline (SAL)
+Developing a weakly-supervised classification pipeline for discharge letters requires a procedure to derive weak-labels for at least a subset of data.
+Weak-labels are labels that can be automatically associated to the data but that might contain some noise.
+This section presents a methodology to construct these weak labels by extracting diagnosis strings from discharge letters and clustering them.
 This allows the creation of groups of diagnoses related to the same disease or symptoms that can be exploited as labels.
 This step corresponds to Step 1 (SAL) of Figure 1.
-Our approach combines regular expressions for extracting diagnosis strings and a customized clustering pipeline based on the BERT model.
-In the following paragraphs, we describe in detail these sub-steps.
-Extraction of diagnosis strings
-The first step, corresponding to block 1.1 in Figure 1, is the extraction of diagnosis strings.
+Our approach combines regular expressions for extracting diagnosis strings and a customized clustering pipeline based on a transformer model.
+In the following sections, we describe in detail these sub-steps.
+2.2.1 Extraction of diagnosis strings
+The first step (1.1 in Figure 1.a) is the extraction of diagnosis strings.
 Italian discharge letters do not follow a standardized template, as each hospital has its own, and even different departments within the same hospital may personalize the template further.
@@
 Regular expressions provide a powerful and flexible approach for pattern matching, making them suitable for capturing a template field with limited variations, such as the diagnosis indication in our dataset.
-We reported in Supplementary Material C the expressions we defined to extract the diagnosis strings.
+C reports the set of regular expressions used to extract the diagnosis strings.
 After this step, we have, for the majority of the letters, a string, typically a short sentence, explaining the reasons for the patient’s admission.
@@
 To do this, it is necessary to cluster them, as detailed in the following paragraph.
-Diagnosis Clustering
+2.2.2 Diagnosis Clustering
 Once the diagnosis strings are extracted, we use a tailored clustering pipeline to group them meaningfully, aiming at identifying clusters related to the various types of diseases and symptoms.
-This step corresponds to block 1.2 in Figure 1 and is expanded in Figure 2.
-Figure 2: Clustering section of the pipeline. Detailed diagram of the clustering pipeline for diagnosis strings, corresponding to Step 1.2 of the full pipeline depicted in Figure 1.
-Punctuation removal, digits removal and abbreviation expansion are applied to the diagnosis strings (step 1.2.1).
-This pre-processing aims to facilitate the subsequent construction of a text representation that can be representative of the semantics of the text.
-In particular, punctuation signs and digits are irrelevant for discriminating between diagnoses and can be safely removed.
-The problem of abbreviations is a common problem in medical documents: a wide variety of acronyms exist in the medical domain, and their meaning can also be context-dependent.
-We have extracted all the words with two and three letters from our dataset of diagnosis strings and expanded those that we found to be medical abbreviations.
+This corresponds to Step 1.2 in panel a of Figure 1 and is expanded in panel b of the same figure.
+Minor pre-processing steps, including the removal of punctuation, dates, and time expressions, are applied to facilitate the subsequent construction of a text representation that captures the semantic content of the document, while excluding elements that are not informative for discriminating among diagnoses (Step 1.2.1).
 After this pre-processing, the pipeline needs to construct an embedding of the text that a clustering algorithm will subsequently use.
-This step (1.2.2) exploits the BERT model we discussed in Section 2.0.1 (BPT).
-The BERT-based models produce embedding vectors whose size is 768, making it necessary to apply a dimensionality reduction technique over these embeddings to use them as input for a clustering algorithm (Step 1.2.3).
+This step (1.2.2) exploits the transformer model we discussed in Section 2.1.
+The transformer-based model produces embedding vectors whose size is 768, making it necessary to apply a dimensionality reduction technique (PCA) over these embeddings to use them as input for a clustering algorithm (Step 1.2.3).
 After this step, a clustering algorithm can be applied to the resulting embedding vectors (Step 1.2.4).
-We selected HDBSCAN since it provides many relevant advantages: it does not require defining a priori the number of clusters, it does not necessarily cluster all the data, and it can identify clusters of varying shapes and sizes.
+We selected HDBSCAN [campello2013density] because it does not require specifying the number of clusters a priori, allows leaving part of data unclustered, and can identify clusters with heterogeneous shapes and densities.
+These properties are well suited to transfromer-based text embeddings, which often lie on complex, nonlinear manifolds and may form clusters with varying densities, irregular geometries, and overlapping boundaries.
+Several commonly used clustering algorithms, such as K-Means, Gaussian Mixture Models, and standard hierarchical clustering, are less suitable for this type of data [asyaky2021improving, yeasmin2022transformer], as further supported by a comparative analysis reported in D.
 To effectively group strings related to different diseases and symptoms, the clustering must prioritize the intra-cluster similarity against the inter-cluster dissimilarity.
@@
 We employ a keyword representation in which, for each cluster, we extract words that appear with significant frequency in the cluster and do not appear with considerable frequency over the entire dataset (Step 1.2.5).
-These keywords can be represented with BERT embeddings in the same way as the original diagnosis strings and used for subsequent clustering (Step 1.2.6).
-Mapping clusters to diseases
+These keywords can be represented with transformer embeddings in the same way as the original diagnosis strings and used for subsequent clustering (Step 1.2.6).
+2.2.3 Mapping clusters to diseases
 Step 1.3 consists of mapping the clusters to the weak labels that will be used by the classification model.
@@
 Similarly, manually reviewing the original diagnosis strings is impractical due to their large number.
-2.0.3 Classification of discharge letters using the weakly-labelled dataset (WLC)
-The classification pipeline that exploits the weakly labelled data produced by Step 1 corresponds to Step 2 in Figure 1 (WLC).
-A pre-processing step (Step 2.1) is necessary to remove headers and footers that contain information that is non-relevant to our analysis.
-Supplementary Material B reports details about this.
-The classification model we adopted is based on the fine-tuned version of the Umberto model that we presented in Section 2.0.1 (BPT).
+2.3 Classification of discharge letters using the weakly-labelled dataset (WLC)
+The classification pipeline that exploits the weakly-labelled data produced by Step 1 corresponds to Step 2 (WLC) in Figure 1.a.
+A pre-processing step (Step 2.1) removes headers and footers that contain information that is non-relevant to our analysis (see B for details and results without pre-processing).
+The classification model we adopted is based on the fine-tuned version of the UmBERTo model that we presented in Section 2.1.
 In this case, the transformer-based model is enhanced with a linear classification layer that outputs the probability of having the disease.
 It is trained with the discharge letters’ text and the weak labels assigned to the letters based on the clusters to which their diagnoses have been assigned (Step 2.2).
-Given the limitation of BERT-based models in dealing with texts longer than 512 tokens, when this is the case, we truncated the letter text to the first 512 tokens after pre-processing, possibly including the diagnosis sentences, even if at the end of the letter (in the setting in which we kept it).
+Given the limitation of RoBERTa-based models in dealing with texts longer than 512 tokens, when this is the case, we truncated the letter text to the first 512 tokens after pre-processing, possibly including the diagnosis sentences, even if at the end of the letter (in the setting in which we kept it).
 This is not a relevant limitation, considering that 90% of discharge letters in our dataset are within 512 tokens after pre-processing.
 In the case of datasets with longer texts, it is possible to chunk them in portions no longer than 512 tokens, process them in parallel and combine their outputs, taking the maximum probability.
-2.1 Training and evaluation
-The BERT-based classifier is trained with the AdamW optimizer, using a linear learning rate scheduler starting from $10^{-3}$, with a weight decay of 0.01, for six epochs.
-Initially, only the last classification layer is trained, and after two epochs, the last BERT layer is unfrozen.
-After four epochs, the last but one BERT layer is unfrozen.
-These hyperparameters have been selected with grid search.
-We cannot perform a precise separate evaluation of all the pipeline steps since we do not have ground truth labels related to diagnosis strings and all their clusters.
-Because of this, we assess the performance of the whole pipeline on the classification task, implicitly evaluating all the previous steps.
-For the clustering, we can perform a specific evaluation limited to the clusters related to our bronchiolitis gold labels.
-[…]
-[…]
-[…]
-The classification results are evaluated in terms of precision, recall and F1-score for the positive class and area under the receiver operating characteristic curve (AUC).
-These metrics are computed for both our weak labels and the gold standard bronchiolitis labels.
-Evaluating against weak labels allows for an assessment of the classification model’s performance alone, while evaluation against gold standard labels provides a comprehensive view of the pipeline’s effectiveness.
-We exclude AUC for rule-based classifiers, as they do not produce probability estimates, and we do not assess the fully supervised classifier using weak labels.
-We compute the metrics with 10-fold stratified cross-validation.
+2.4 Experimental outline
+Our main evaluation is at the final classification stage of the pipeline, where we can compare the results against manually annotated labels.
+In addition to this, we evaluated the intermediate steps of the pipeline, in particular the diagnosis extraction (Step 1.1) and clustering (Steps 1.2-1.3), on a randomly selected subset of 1,000 diagnosis strings that have been manually annotated.
+Each of these strings was marked as correct, partially correct , or wrong with respect to the fact of being a complete diagnosis string and to the assignments to the first level and the second level clusters.
+Additional results on the clustering steps, including robustness analyses with respect to hyperparameter choices and clustering algorithms, are reported in D.
+In evaluating classification results, we compare our pipeline with some fully unsupervised baselines: two rule-based approaches that do not require supervision and with a zero-shot LLM for the Italian language (LLaMAntino-3-ANITA-8B-Inst-DPO-ITA[polignano2024advanced].
+The two rule-based approaches look for the presence of the terms corresponding to the definitions of bronchiolitis used in Step 1.3, with the first one (Rule-based full-text) looking into the entire text of the letter and the second one (Rule-based diagnosis) looking only in the diagnosis strings extracted by Step 1.1.
+The zero-shot LLM is prompted to indicate if the text of the text discharge letter indicates a diagnosis of bronchiolitis, answering with a “Si” (“Yes”) or a “No”.
+Additionally, we assess in detail the performance of the classification model of our pipeline (Step 2.2) by replacing it with other classifiers:
+  1. 1.
+    The original UmBERTo model [tamburini2020bertology], without TPT (WS-UmBERTo)
+  2. 2.
+    XLM-RoBERTa-base [conneau2020unsupervised], a multilingual version of RoBERTa [liu2019roberta], with (WS-XLM-RoBERTa-TPT) and without (WS-XLM-RoBERTa) TPT
+  3. 3.
+    RemBERT [chung2020rethinking], an improved version of the multilingual mBERT, with (WS-RemBERT-TPT) and without (WS-RemBERT-BPT) without TPT
+  4. 4.
+    An LSTM-based recurrent neural network exploiting Italian Word2Vec embeddings [mikolov2013efficient] (WS-LSTM)
+The three transformer-based models with pretraining are also evaluated in a fully supervised setting, training them on the manually annotated labels (FS-UmBERTo-TPT, FS-XLM-RoBERTa-TPT, FS-RemBERT-TPT), so as to have a measure of the performance loss due to weak supervision.
 Considering that our weak labels are derived from clusters of diagnosis strings that we extracted from discharge letters, this might bias the classification models into looking only into the diagnosis sections of the letters.
@@
   2. 2.
-    Removing the diagnosis strings exploited to assign weak labels
-Since our dataset contains data from different hospitals and different LHUs, we verify the presence of differences among them.
-Since the number of cases for each hospital and LHU is limited, we consider those with a minimum number of bronchiolitis cases (15) and measure the results using a Leave One Group Out (LOGO) approach.
-For each hospital/LHU to be evaluated, we train the model on the data from the other hospitals/LHUs, and we measure the performance on the data from the hospital/LHU of interest.
-We also compare results between the groups of pediatric vs non-pediatric ER/Department admissions.
-In this case, we adopt the same 10-fold stratified cross-validation used to compare different classification models, dividing the test set between pediatric and non-pediatric ERs/Departments at each iteration and computing the metrics on these two groups separately.
+    Removing the diagnosis strings exploited to assign weak labels (when present)
+We also measure the performance of the models on the weak labels, to assess the difference with the ones observed on gold labels, and we verify the robustness of the results on gold labels to the weak label selection, measuring the effect of removing single clusters from positive weak labels.
+To further analyze the noise introduced by weak labels, we also experiment with replacing a varying percentage of gold labels with weak labels.
+The classification results are evaluated in terms of precision (P), recall (R) and F1-score (F1) for the bronchiolitis class and area under the receiver operating characteristic curve (AUC).
+For classifiers producing a continuous score, the classification threshold is set via Youden’s index [youden1950index].
+We exclude AUC for rule-based classifiers, as they do not produce probability estimates, and we do not assess the fully supervised models using weak labels.
+The positive class probability for the zero-shot LLM is estimated by applying a softmax function over the token-level log-likelihoods of the tokens “Si” (“Yes”) and “No”, following previous literature [brown2020language].
+All metrics are computed with 10-fold stratified cross-validation.
+Since our data come from different hospitals and different LHUs, we also perform additional experiments, reported in E, to verify the presence of differences among them.
+2.5 Training details
+The additional pre-training of the transformer-based models over the E3C corpus was performed with the masked-word prediction task, with a learning rate of $2\cdot 10^{-5}$ and weight decay of $0.01$, using 20% of the data as a validation set for early stopping.
+The LSTM baseline has the frozen Word2Vec embedding layer followed by a 32-unit LSTM layer, a dropout layer, a 16-unit fully connected layer, a dropout layer and the final output layer.
+It is trained with Adam optimizer, with a learning rate scheduler starting from $10^{-2}$, a decay rate of 0.7, 100 decay steps, using early stopping.
+The transformer-based classifiers are trained with the AdamW optimizer, using a linear learning rate scheduler starting from $10^{-3}$, with a weight decay of 0.01, for a number of epochs regulated by early stopping (leading to 6 epochs for the UmBERTo models, 4 for the XLM-RoBERTa models and 3 for the RemBERT models).
+Initially, only the last classification layer is trained, and after two epochs, the last transformer layer is unfrozen.
+After four epochs, if reached, the last but one transformer layer is unfrozen.
+These hyperparameters have been selected with grid search.
+Due to the class imbalance, in the training folds the negative class is undersampled, limiting its size to twice that of the positive class.
 3 Results
-In this section, we present the results of our pipeline for the bronchiolitis case study, for the different stages that compose it.
-In particular, Section 3.1 describes the dataset, Section 3.2 presents results related to the extraction of diagnosis strings from discharge letters, Section 3.3, covers the clustering pipeline and Section 3.4 presents the final results on the identification of bronchiolitis.
 3.1 Dataset
-The dataset comprises 33’176 discharge letters related to children who were admitted to Emergency Rooms and/or hospitalized in the Veneto Region between 2017 and 2020.
+The dataset for the case study comprises 33,176 discharge letters related to children who were admitted to Emergency Rooms and/or hospitalized in the Veneto Region between 2017 and 2020, with a family pediatrician adhering to the Pedianet network [cantarutti2021pedianet].
+Inclusion in the Pedianet database is voluntary and parents/legal guardians must provide informed consent for their children’s anonymized data to be used for research purposes.
+The study was performed in accordance with the Declaration of Helsinki.
+Ethical approval of the study and access to the database were granted by the Internal Scientific Committee of So.Se.Te.
+Srl, the legal owner of Pedianet.
 The dataset covers 44 different hospitals in the region, which can be grouped into 9 Local Health Units (LHUs) and 2 university hospitals.
-Each discharge letter in the dataset includes the complete text of the letter, along with the date of the letter and anonymized identifiers of the patient.
+Each discharge letter includes the complete text of the letter, along with the date of the letter and an anonymized identifier of the patient.
 A manual review of discharge letters was conducted to extract relevant diagnoses of bronchiolitis, resulting in 335 positive cases within the dataset.
 This review was performed by a data manager with a clinical background, with the support of clinicians for uncertain cases.
-These labels are the gold standard to validate our pipeline, and we do not make direct use of them in the pipeline training.
-We identified the records related to pediatric ERs/Departments by looking at the presence of specific keywords (e.g., pediatria, pediatrico, pediatrica) in the header of the letters.
-This marked 10’340 records (31.17%) as related to pediatric ERs/Departments.
-Supplementary Material A1 reports additional plots with data distributions among hospitals and LHUs.
-In particular, ten hospitals cover more than 85% of the records.
-The median length of the letters is 4515 characters (1st-3rd quartiles 3240-5806). An example of a fictitious letter is reported in Supplementary Material A2.
+These labels are the gold standard to validate our pipeline, and we do not make direct use of them in the pipeline training (except for the fully-supervised experiments).
+A.1 reports additional plots with data distributions among hospitals and LHUs. In particular, 10 hospitals cover more than 85% of the records.
+The median length of the letters is 4,515 characters (1st-3rd quartiles 3,240-5,806). An example of a fictitious letter is reported in A.2.
 3.2 Diagnosis extraction
 Our diagnosis extraction module (Step 1.1) successfully extracted a diagnosis sentence from 89% of the discharge letters in the dataset.
-We extracted 15’038 unique diagnosis sentences, of which 75% occur only once, confirming the need for clustering them in wider groups that can be more easily related to known diseases and symptoms.
-A few of the extracted strings include the ICD9-CM codes and their standard textual description.
-Specifically, 737 diagnosis strings (4.9%) contain an ICD9-CM code, covering 3’436 discharge letters (10.4%).
+We extracted 15,038 unique diagnosis sentences, of which 75% occurred only once, confirming the need for clustering them in wider groups that can be more easily related to known diseases and symptoms.
+ICD-9-CM codes are present only in 737 diagnosis strings (4.9%), covering 3,436 discharge letters (10.4%).
 The most common strings are the simplest, while more articulate ones have single occurrences.
-The median length of the extracted diagnoses strings is 49 characters, with 1st and 3rd quartiles of 17 and 156 characters.
+The median length of the extracted diagnoses strings is 49 characters, with first and third quartiles of 17 and 156 characters.
 Among the 335 bronchiolitis-labelled letters, we extracted 138 unique diagnosis strings from 306 letters.
+The manual validation of a sample of 1,000 extracted diagnosis strings revealed that 985 (98.5%) were correct diagnosis strings, only 4 (0.4%) were not diagnosis strings, and 11 (1.1%) were incomplete diagnosis strings.
 3.3 Diagnosis clustering
-In the first clustering level, the 15’038 extracted diagnosis strings were grouped into 590 clusters.
+In the first clustering level, the 15,038 extracted diagnosis strings were grouped into 590 clusters.
 Cluster sizes varied significantly, from a large cluster of 270 strings to several smaller clusters containing 5 strings.
-Table 1 presents the 20 largest clusters, including their sizes and descriptive keywords.
 The subsequent second-level clustering reduced the number of clusters to 135.
+Summary of main clusters is reported in D.
+Table 1: Manual evaluation of the clustering assignment for 1,000 diagnosis strings. comparing results after first level and second level clustering.
+Correct Partially Correct Wrong
+1st level clustering 69.60 % 22.20 % 8.20 %
+2nd level clustering 71.90 % 20.30 % 7.80 %
+$\Delta$ 2nd-1st 2.30 % -1.90 % -0.40 %
+Table 1 summarizes the manual evaluation of the clustering assignment, at both levels, for the annotated sample of 1,000 strings.
+Most of the partially correct cluster assignments are due to details of the diseases, e.g., fracture of right hand assigned to a cluster for fracture of left hand, severe influenza assigned to a cluster for mild influenza.
+A few cases are related to multiple symptoms/diseases not completely matching the cluster, such as vomiting and fever in the cluster for fever alone.
 For our case study on bronchiolitis identification, Table 2 details the main clusters related to respiratory infections, including the number of bronchiolitis cases they cover.
+Given space constraints, we report here only clusters mentioning bronchiolitis, bronchospasm, pneumonia, or respiratory infections, as these are the clusters most closely related to bronchiolitis or its associated symptoms.
 Four clusters (#1, #2, #3, #5) contain the keyword “bronchiolitis”.
-Three include a large majority of bronchiolitis cases, whereas the fourth cluster (#5) has only half of its diagnoses labelled as bronchiolitis.
+Three of them include a large majority of bronchiolitis cases, whereas the fourth cluster (#5) has only half of its diagnoses labelled as bronchiolitis.
 These four clusters together cover 65.38% of all bronchiolitis cases.
-Additionally, 13 clusters include the keyword “bronchospasm”.
+Additionally, 13 clusters include the keyword “bronchospasm”, a symptom frequently associated with bronchiolitis.
 According to physicians who annotated the data, bronchospasm is associated with bronchiolitis only if accompanied by fever.
 Only one of these 13 clusters (#4) has both “bronchospasm” and “fever” keywords, and indeed, it contains a high percentage (70%) of bronchiolitis, accounting for an additional 15% of total bronchiolitis cases.
-Thus, these five clusters are those that, without knowing the gold labels, would be selected to define the weak labels for the training set, according to the two definitions of {bronchiolitis} and {bronchospasm, fever}.
+Thus, these five clusters are those that, without knowing the gold labels, would be selected to define the weak labels for the training set, according to the two keyword-based definitions of {positive={bronchiolitis},negative={}} and {positive={bronchospasm, fever},negative={}}.
 Among the respiratory infection clusters, three other clusters contain bronchiolitis cases (#6, #7, #8), but they are in a limited number.
 For the second-level clustering, Table 3 presents clusters related to respiratory infections.
-Some clusters from Table 2 have been merged: two related to mild bronchiolitis, eight related to bronchospasm, and five about generic respiratory infections.
+The second level clustering appears to have effectively merged some of the clusters that were previously separated in Table 2: two related to mild bronchiolitis, eight related to bronchospasm, and five about generic respiratory infections.
 The selection of weak labels for training remains unchanged since, looking at their keywords, we select second-level clusters #1, #2, #3, #4 that cover the same records as the clusters we would have chosen at the first level.
-The main advantage of second-level clustering is the substantial reduction in the number of clusters (approximately 1/5 of the original number of clusters in our dataset), simplifying weak labelling, especially if manual verification is involved.
+Second-level clustering determines a substantial reduction in the number of clusters (approximately 1/5 of the original number of clusters in our dataset), simplifying manual inspection and determining a positive effect on diagnosis-cluster mappings as observed in Table 1.
 This cluster selection for weak labels covers 80% of the original bronchiolitis cases.
-To understand the reason for the false negatives, we reviewed twenty letters labelled as bronchiolitis in the gold standard but not covered by our weak labels.
+To understand the reason for the false negatives, we reviewed 20 letters labelled as bronchiolitis in the gold standard but not covered by our weak labels.
 We found 10 cases where the diagnosis pertains to other symptoms or disease (e.g., otitis, dyspnea, fever), but the bronchiolitis is mentioned in the letter, at least as a suspect condition; 4 cases where bronchiolitis is mentioned in the diagnosis alongside other diseases, leading to assignment to a cluster related to the other disease; 3 cases of respiratory issues with no explicit mention of bronchiolitis, respiratory syncytial virus, or bronchospasm and fever; 2 cases with no diagnosis string present; 1 case where a family member has bronchiolitis, but the patient’s diagnosis is of a different disease.
 These results suggest that the absence of some gold labels in our weak labels primarily arises from the lack of explicit bronchiolitis indications, with only a few cases that could have been detected but were missed.
-This highlights the importance of using the full discharge letters for a model aiming to identify diseases accurately.
-Furthermore, there might be a few errors in the gold labels themselves.
-We cannot perform a complete evaluation of the clusters produced using our Umberto-E3C model against those produced using the original Umberto model, but we can compare the distribution of gold bronchiolitis labels on the two sets of clusters.
-Supplementary Material D reports the clusters obtained using the original Umberto model in our clustering pipeline.
-The number of clusters related to respiratory infections is higher (34 vs 27), and their keywords appear more heterogeneous.
+They also highlight the importance of using the full discharge letters for a model aiming to identify diseases accurately.
+To assess the effect of the TPT step on clustering results, D reports the clusters obtained using the original UmBERTo model in our clustering pipeline.
+The number of clusters related to respiratory infections is higher without TPT (34 vs 27), and their keywords appear more heterogeneous.
 Selecting weak labels by applying the same criterion we used on our clusters, we would cover only 47% of bronchiolitis cases with the clusters from the original UmBERTo model.
-Moreover, they are less compact, introducing more false negatives in the weak labels.
-Table 1: Keyword-based representation of the twenty largest first-level diagnosis clusters
+Moreover, these clusters are more hetereogenous since they introduce more false negatives in the weak labels.
+Additional results on the effect of clustering algorithms and hyperparameters are reported in D.
+Table 2: Keyword-based representation of the first-level clusters related to respiratory infections, with the % of bronchiolitis cases in each cluster and the % to the total number of cases in the dataset
+Cluster # Cluster keywords [ITA] Cluster keywords [ENG]
+Within cluster
+bronch. %
+% of total bronch.
+covered
+1 bronchiolite lieve mild bronchiolitis 85.19 6.87
+2 acuta bronchiolite iniziale lieve acute mild initial bronchiolitis 85.14 44.48
+3 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 74.05 11.94
+4 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 70.00 15.02
+5 alimentazione bronchiolite difficolta’ feeding difficulties bronchiolitis 50.00 2.09
+6 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 12.70 2.39
+7 broncospasmo infezione respiratoria bronchospasm respiratory infection 6.67 0.30
+8 broncospasmo disidratazione lieve bronchospasm mild dehydration 1.17 0.60
+9 flogosi alte vie respiratorie febbre high airways flogosis fever 0.00 0.00
+10 broncospasmo infezione vie respiratorie bronchospasm airways infection 0.00 0.00
+11 acuto broncospasmo acute bronchospasm 0.00 0.00
+12 broncospasmo flogosi vie bronchospasm airways flogosis 0.00 0.00
+13 broncospasmo polmonite bronchospasm pneumonia 0.00 0.00
+14 broncospasmo corso ongoing bronchospasm 0.00 0.00
+15 broncopolmonite sinistra left bronchopneumonia 0.00 0.00
+16 febbricola febbrile infezione respiratoria virosi febrile fever respiratory infection virosis 0.00 0.00
+17 broncospasmo episodio bronchospasm episode 0.00 0.00
+18 infezione alte vie respiratorie high airways infection 0.00 0.00
+19 flogosi infezione alte vie respiratorie flogosis high airways infection 0.00 0.00
+20 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis 0.00 0.00
+21 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm 0.00 0.00
+22 broncospasmo otite bronchospasm otitis 0.00 0.00
+23 broncospasmo parainfettivo bronchospasm parainfective 0.00 0.00
+24 broncospasmo virosi bronchospasm virosis 0.00 0.00
+25 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory 0.00 0.00
+26 crisi epilessia infezione paziente polmonare respiratoria seizure epilepsy respiratory infection lung patient 0.00 0.00
+27 altre infezioni acute vie respiratorie other acute respiratory infection airways 0.00 0.00
+Table 3: Keyword-based representation of the second-level clusters related to respiratory infections, with the % of bronchiolitis cases in each cluster and the % to the total number of cases in the dataset
+2nd level cluster # 2nd level clusters keywords [ITA] 2nd level clusters keywords [ENG] 1st level cluster # 1st level clusters keywords [ITA] 1st level clusters keywords [ENG] Within cluster bronch.
+% % of total bronch.
+covered 1 bronchiolite lieve mild bronchiolitis 1 acuta bronchiolite iniziale lieve acute mild initial bronchiolitis 85.16 51.34 2 bronchiolite lieve mild bronchiolitis 2 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 3 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 74.05 11.94 3 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 4 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 70.00 15.02 4 bronchiolite difficolta di alimentazione feeding difficulties bronchiolitis 5 bronchiolite difficolta di alimentazione feeding difficulties bronchiolitis 50.00 2.09 5 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 6 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 12.7 2.39 6 broncospasmo infezione respiratoria bronchospasm respiratory infection 7 broncospasmo infezione vie respiratorie bronchospasm infection airways 2.9 0.30 10 broncospasmo infezione respiratoria bronchospasm respiratory infection 7 broncospasmo disidratazione lieve bronchospasm mild dehydration 8 broncospasmo disidratazione lieve bronchospasm mild dehydration 1.17 0.60 8 broncospasmo bronchospasm 11 acuto broncospasmo acute bronchospasm 0.00 0.00 12 broncospasmo flogosi vie bronchospasm airways flogosis 13 broncospasmo polmonite bronchospasm pneumonia 14 broncospasmo corso ongoing bronchospasm 22 broncospasmo otite bronchospasm otitis 23 broncospasmo parainfettivo bronchospasm parainfective 24 broncospasmo virosi bronchospasm virosis 17 broncospasmo episodio bronchospasm episode 9 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis 20 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis 0.00 0.00 10 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm 21 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm 0.00 0.00 11 infezione vie respiratorie airways infection 18 infezione alte vie respiratorie high airways infection 0.00 0.00 19 flogosi infezione alte vie respiratorie flogosis high airways infection 0.00 0.00 27 altre infezioni acute vie respiratorie other acute respiratory infection airways 0.00 0.00 26 crisi epilessia infezione paziente polmonare respiratoria seizure epilepsy respiratory infection lung patient 0.00 0.00 16 febbricola febbrile infezione respiratoria virosi febrile fever respiratory infection virosis 0.00 0.00 12 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory 25 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory 0.00 0.00 13 alte febbre flogosi respiratorie vie high fever flogosis airways 9 flogosi alte vie respiratorie febbre high airways flogosis fever 0.00 0.00 14 broncopolmonite sinistra left bronchopneumonia 15 broncopolmonite sinistra left bronchopneumonia 0.00 0.00
+3.4 Disease classification
+Table 4: 10-fold CV results for classification on gold labels. Std. Dev. in parenthesis.
+Learning Classifier P [%] R [%] F1 [%] AUC [%]
+With diagnosis string
+Weakly-supervised WS-UmBERTo-TPT
+78.34
+77.57
+78.14
+77.68
+WS-XLM-RoBERTa-TPT
+69.54
+71.33
+70.85
+69.24
+WS-RemBERT-TPT
+67.41
+70.11
+69.23
+68.00
+WS-UmBERTo
+68.29
+75.99
+72.15
+71.21
+WS-XLM-RoBERTa
+67.49
+68.67
+68.02
+67.65
+WS-RemBERT
+65.24
+63.82
+64.96
+65.74
+WS-LSTM
+57.39
+60.86
+59.87
+63.29
+Unsupervised Rule-based full-text
+65.57
+68.76
+67.22
+Rule-based diagnosis
+74.61
+67.42
+70.74
+LLaMAntino-3
+78.26
+74.64
+75.12
+71.67
+Supervised FS-UmBERTo-TPT
+80.30
+81.52
+80.83
+83.45
+FS-XLM-RoBERTa-TPT
+75.29
+78.31
+76.54
+78.54
+FS-RemBERT-TPT
+73.24
+75.28
+74.21
+76.39
+Without diagnosis string
+Weakly-supervised WS-UmBERTo-TPT
+74.25
+73.52
+73.89
+74.27
+WS-XLM-RoBERTa-TPT
+66.83
+67.52
+67.19
+67.25
+WS-RemBERT-TPT
+64.53
+63.66
+64.21
+65.47
+WS-UmBERTo
+67.55
+76.40
+71.74
+69.98
+WS-XLM-RoBERTa
+63.98
+65.04
+64.69
+64.87
+WS-RemBERT
+62.25
+60.48
+61.65
+62.21
+WS-LSTM
+50.14
+63.97
+56.34
+59.02
+Unsupervised Rule-based full-text
+62.33
+60.05
+61.15
+LLaMAntino-3
+70.42
+72.98
+71.16
+68.52
+Supervised FS-UmBERTo-TPT
+78.25
+76.21
+77.15
+78.51
+FS-XLM-RoBERTa-TPT
+72.54
+73.65
+72.96
+74.08
+FS-RemBERT-TPT
+71.49
+70.87
+71.21
+72.68
+Table 5: 10-fold CV results for classification on weak labels. Std. Dev. in parenthesis.
+Learning Classifier P [%] R [%] F1 [%] AUC [%]
+With diagnosis string
+Weakly-supervised WS-UmBERTo-TPT
+82.25
+80.24
+81.35
+82.30
+WS-XLM-RoBERTa-TPT
+76.74
+74.91
+75.89
+75.53
+WS-RemBERT-TPT
+76.34
+75.57
+75.85
+74.51
+WS-UmBERTo
+80.67
+75.48
+78.01
+78.67
+WS-XLM-RoBERTa
+75.53
+70.73
+73.24
+73.82
+WS-RemBERT
+74.15
+65.35
+70.41
+72.51
+WS-LSTM
+68.57
+64.97
+65.52
+67.50
+Unsupervised Rule-based full-text
+70.18
+87.05
+77.74
+Rule-based diagnosis
+73.88
+86.38
+79.85
+LLaMAntino-3
+57.39
+63.78
+79.64
+70.87
+Without diagnosis string
+Weakly-supervised WS-UmBERTo-TPT
+84.43
+77.63
+80.79
+80.57
+WS-XLM-RoBERTa-TPT
+74.47
+71.68
+73.11
+75.62
+WS-RemBERT-TPT
+72.13
+72.64
+72.40
+72.27
+WS-UmBERTo
+80.08
+72.15
+76.02
+76.16
+WS-XLM-RoBERTa
+72.21
+65.56
+68.74
+69.62
+WS-RemBERT
+69.86
+68.32
+69.25
+70.34
+WS-LSTM
+64.98
+56.22
+60.54
+63.00
+Unsupervised Rule-based full-text
+65.54
+79.16
+71.28
+LLaMAntino-3
+49.81
+79.91
+61.12
+65.32
+Table 4 presents the 10-fold cross-validation results for the classification task, comparing our weakly-supervised pipeline in versions with different transformer-based classifiers, fully unsupervised rule-based and zero-shot LLM approaches, and fully-supervised versions of the transformer-based classifiers.
+Results are reported for both the setting where we keep the diagnosis string and the one where we remove it from the letter.
+The same results are reported in Table 5 evaluating on the weak labels.
+Our weakly-supervised pipeline outperform fully unsupervised approaches and maintains a limited gap with the fully supervised models.
+The TPT step consistently improved performance across models: the differences are limited, but there are no cases in which TPT hurts the performance.
+Among the transformer-based classifiers, UmBERTo constantly outperforms RemBERT and XLM-RoBERTa.
+The LSTM classifier has a significant performance gap with the transformer-based ones.
+Rule-based full-text, which verifies the keywords in the entire letter, shows lower precision and recall than the majority of the weakly-supervised models, emphasizing the need for a more sophisticated approach.
+The results further decline when the diagnosis string is excluded.
+Rule-based diagnosis, using only diagnosis strings, yields better precision than the rule applied to the complete text, though recall remains similar.
+The zero-shot approach with the Italian LLaMAntino-3 LLM does not outperform our best weakly-supervised results, but shows competitive results.
+When comparing the classification results between weak and gold labels, we observe a decrease, particularly in precision.
+Notably, LLaMAntino-3 is the only model which shows better results on gold labels than on weak labels.
+While this might be expected for models that were not trained on weak labels, this does not happen with rule-based methods that are probably more aligned to the way in which weak labels are identified.
+The transformers-based models show limited differences between settings with and without diagnosis string, while these are larger for the LSTM classifier.
+As expected, the fully supervised models demonstrated better overall performance, though differences are modest.
+Figure 2: AUC with varying percentages of weak labels.
+Mean AUC ($\pm 95\%$ CI) on gold labels, with and without diagnosis strings, as different proportions of gold labels are replaced by weak labels.
+Results are averaged over 5 repetitions of 10-fold cross-validation.
+In Figure 2, we observe the decrease in performance when we replace increasing percentages of gold labels with weak labels.
+This picture is representative of many real world settings where only a certain percentage of the data can have gold labels, and shows a smooth and monotonic decrease.
+In Table 6, we assess the differences in the results of our best weakly-supervised model (WS-UmBERTo-TPT) when selecting different subsets of clusters for the weak labels.
+In particular, we exclude one of the 1st level clusters we selected at a time.
+Excluding clusters #5 or #1 does not affect the results, which is reasonable considering their smaller size.
+Removing clusters #3 or #4 has a negative effect, but limited.
+Only excluding cluster #2 determines a drop in the performance and an increase in the variance, which is expected since it accounts for almost half of the bronchiolitis cases.
+These results confirm the robustness of the model with respect to small changes in weak labels selection.
+E reports additional results comparing different hospitals, LHUs and pediatric vs non-pediatric ER/Departments.
+Some hospitals are more represented in the dataset, leading to a superior performance, but also hospitals with a similar number of records show considerable performance variability, suggesting that reporting practices in discharge letters might influence disease detection.
+Table 6: 10-fold CV results for classification, considering different groups of 1st level clusters for weak labels. Std. Dev. in parenthesis. G = gold labels, W = weak labels
+Clusters for weak labels
+(1st level)
+P [%] (W)
+R [%](W)
+F1 [%] (W)
+AUC [%] (W)
+P [%] (G)
+R [%] (G)
+F1 [%] (G)
+AUC [%] (G)
+#1. #2. #3. #4 78.39 (7.04) 73.35 (6.21) 75.79 (4.65) 83.02 (4.16) 70.75 (6.48) 81.27 (6.99) 75.65 (4.29) 75.49 (5.57)
+#1. #2. #4. #5 73.69 (8.89) 83.20 (9.16) 78.16 (7.80) 80.71 (7.01) 68.42 (8.70) 75.09 (8.58) 71.60 (7.26) 72.67 (6.87)
+#2. #3. #4. #5 83.49 (7.87) 74.66 (6.88) 78.82 (4.26) 84.02 (4.43) 68.54 (6.30) 82.35 (4.94) 74.81 (3.82) 75.83 (5.22)
+#1. #3. #4. #5 93.73 (6.85) 78.11 (8.47) 85.19 (7.06) 90.91 (7.35) 57.29 (12.97) 59.03 (15.94) 58.15 (12.45) 68.90 (10.04)
+#1. #2. #3. #5 80.45 (8.90) 74.62 (8.57) 77.43 (5.76) 85.95 (5.29) 67.86 (6.36) 73.42 (8.64) 70.53 (5.33) 70.22 (3.10)
+4 Discussion
+This study introduces a novel pipeline demonstrating how weakly-supervised learning can be effectively applied to clinical NLP.
+To our knowledge, this is the first work addressing diagnosis identification from Italian discharge letters.
+Although we validated it using a case study on bronchiolitis, the pipeline is designed to be applicable to any disease.
+Our case study shows excellent results, outperforming existing unsupervised approaches, and underscores the potential for broader application in clinical settings.
+We observed a good mapping between diagnosis strings and their clusters’ descriptions in the manually evaluated subset of strings, and only a limited performance difference between weak and gold labels (4-7 % of AUC), confirming that the SAL step can effectively identify relevant cases as weak labels.
+Despite the inevitable noise that weak labels introduce, this is limited enough to allow the model to perform well when evaluated against gold labels.
+Additionally, the differences between settings where we retained the diagnosis strings and those where we removed them are limited.
+This indicates that the model is effectively learning from the entire text of the discharge letters, making it particularly advantageous when the automatic extraction of diagnosis strings fails.
+Our manual exploration of some weak labels further revealed that diagnosis strings do not always mention the disease responsible for hospitalization.
+In some cases, they only refer to symptoms, and in cases of multiple diseases, not all may be reported.
+These findings highlight the necessity for a classification model based on the complete text of discharge letters, as the one we developed.
+In contrast, rule-based approaches underperformed because diseases like bronchiolitis are not always explicitly mentioned or may appear in negated or hypothetical contexts.
+While more sophisticated rule-based systems could address some of these issues, such approaches lack the scalability and generalizability of our pipeline.
+Conversely, our pipeline can be easily adapted to other diseases, which makes it a versatile tool for clinical NLP.
+The zero shot LLM performed better than rule-based and some weakly-supervised settings, but it did not surpass the best weakly-supervised configuration.
+Despite the reasoning capacities of modern LLMs, their training data rarely include documents and tasks of this type, limiting performance.
+Domain pretraining (TPT) improved results across all models.
+The best performance was obtained with UmBERTo, trained entirely on Italian data, confirming the advantage of language-specific representations even over larger multilingual models.
+Compared with fully supervised classifiers, our weakly supervised pipeline maintained competitive performance, with a limited and acceptable loss ($6\%$ of AUC).
+This is particularly notable given the substantial time and effort required for manual annotation in fully supervised setups: in our case study, annotating the entire dataset would require approximately 3 minutes per discharge letter, amounting to more than 1,500 hours of expert time.
+The experiment in which we progressively replaced gold labels with weak labels (Figure 2) reflects a realistic scenario where only a small portion of the data can be manually annotated.
+Even in this mixed-label setting, performance degraded smoothly and remained close to the fully supervised baseline as long as a modest fraction of gold labels was available, indicating strong resilience to label noise.
+The pipeline has also shown robustness to the cluster selection for the weak labels.
+This highlights its value for settings where large-scale labeling is infeasible.
+While our results are promising, several limitations should be acknowledged.
+First, the gold labels for bronchiolitis were assigned primarily by a single annotator, preventing the estimation of inter-annotator agreement and introducing potential uncertainty in the reference standard.
+Second, the multi-centric nature of our dataset, though a strength in its diversity, also presents an imbalance in the number of records from different hospitals and Local Health Units, which are also all located within a single Italian region.
+This is not only due to the geographical distribution of the population, but might also reflect some biases in the distribution of the family pediatricians adhering to the Pedianet network.
+Testing the pipeline on a more balanced dataset, and externally validating it on data from other regions, would strengthen its generalizability and allow a more precise assessment of performance differences among hospitals.
+Finally, the evaluation of intermediate steps relied on the manual review of a relatively small subset of diagnoses strings.
+Extending this annotation to a larger subset and assessing entire clusters would require substantial effort but could enable a more detailed analysis of the performance of each component in the pipeline.
+Future research can extend this work along multiple directions.
+The pipeline should be validated on multiple diseases, particularly those with overlapping symptoms, using larger annotated datasets.
+Incorporating hospital or LHU identifiers in hierarchical classification models could further improve performance by accounting for site-specific documentation practices.
+Moreover, enhancing the interpretability of model predictions will be crucial for widespread clinical adoption.
+Identifying which portions of each discharge letter most influence the classifier’s decision would help clinicians assess reliability and build trust.
+Techniques from explainable artificial intelligence could be applied to reveal the reasoning process behind model outputs and possibly recognize potential biases present in the training data.
+These future developments would strengthen both the methodological robustness and clinical utility of weakly supervised NLP systems for healthcare text, supporting scalable and trustworthy solutions for real-world applications.
+Declaration of competing interests
+The authors declare that they have no competing interests.
+Data statement
+Data used in this study are not publicly available due to their confidential nature.
+Data are however available upon reasonable request to the corresponding author, subject to approval by the Internal Scientific Committee of Società Servizi Telematici Srl, the legal owner of Pedianet.
+The code developed for the analysis is available at https://github.com/vittot/weakly-supervised-classification-italian-discharge-letters .
+CRediT authorship contribution statement
+Vittorio Torri: Methodology, Software, Investigation, Writing - Original Draft Elisa Barbieri: Conceptualization, Data Curation, Writing - Review & Editing Anna Cantarutti: Conceptualization, Data Curation, Writing - Review & Editing Carlo Giaquinto: Conceptualization, Supervision Francesca Ieva: Conceptualization, Methodology, Writing - Review & Editing, Supervision
+Acknowledgments
+Vittorio Torri and Francesca Ieva have been partially supported by MUR, grant Dipartimento di Eccellenza 2023-2027 awarded to the Department of Mathematics, Politecnico di Milano.
+Elisa Barbieri declares that her position is funded within the project “INF-ACT - One Health Basic and translational Research Actions addressing Unmet Needs on Emerging Infectious Diseases”, ID MUR PE_00000007, under the National Recovery and Resilience Plan (NRRP), Mission 4, Cofmponent 2 – CUP C93C22005170007.
+The authors thank all the family paediatricians collaborating in Pedianet: Eva Alfieri, Michela Alfiero Bordigato, Angelo Alongi, Biagio Amoroso, Rosaria Ancarola, Barbara Andreola, Giampaolo Anese, Roberta Angelini, Maria Grazia Apostolo, Giovanna Argo, Giovanni Avarello, Lucia Azzoni, Maria Carolina Barbazza, Patrizia Barbieri, Gabriele Belluzzi, Eleonora Benetti, Filippo Biasci, Franca Boe, Stefano Bollettini, Francesco Bonaiuto, Anna Maria Bontempelli, Matteo Bonza, Sara Bozzetto, Andrea Bruna, Ivana Brusaterra, Massimo Caccini, Laura Calì, Sonia Camposilvan, Laura Cantalupi, Luigi Cantarutti, Chiara Cardarelli, Giovanna Carli, Sylvia Carnazza, Massimo Castaldo, Stefano Castelli, Monica Cavedagni, Giuseppe Egidio Cera, Chiara Chillemi, Francesca Cichello, Giuseppe Cicione, Carla Ciscato, Mariangela Clerici Schoeller, Samuele Cocchiola, Giuseppe Collacciani, Valeria Conte, Roberta Corro’, Rosaria Costagliola, Nicola Costanzo, Sandra Cozzani, Giancarlo Cuboni, Giorgia Curia, Caterina D’alia, Vito Francesco D’Amanti, Antonio D’Avino, Roberto De Clara, Lorenzo De Giovanni, Annamaria De Marchi, Gigliola Del Ponte, Tiziana Di Giampietro, Giuseppe Di Mauro, Giuseppe Di Santo, Piero Di Saverio, Mattea Dieli, Marco Dolci, Mattia Doria, Dania El Mazloum, Maria Carmen Fadda, Pietro Falco, Mario Fama, Marco Faraci, Maria Immacolata Farina, Tania Favilli, Mariagrazia Federico, Michele Felice, Maurizio Ferraiuolo, Michele Ferretti, Mauro Gabriele Ferretti, Paolo Forcina, Patrizia Foti, Luisa Freo, Ezio Frison, Fabrizio Fusco, Giovanni Gallo, Roberto Gallo, Andrea Galvagno, Alberta Gentili, Pierfrancesco Gentilucci, Giuliana Giampaolo, Francesco Gianfredi, Isabella Giuseppin, Laura Gnesi, Costantino Gobbi, Renza Granzon, Mauro Grelloni, Mirco Grugnetti, Antonina Isca, Urania Elisabetta Lagrasta, Maria Rosaria Letta, Giuseppe Lietti, Cinzia Lista, Ricciardo Lucantonio, Francesco Luise, Enrico Marano, Francesca Marine, Lorenzo Mariniello, Gabriella Marostica, Sergio Masotti, Stefano Meneghetti, Massimo Milani, Stella Vittoria Milone, Donatella Moggia, Angela Maria Monteleone, Pierangela Mussinu, Anna Naccari, Immacolata Naso, Flavia Nicoloso, Cristina Novarini, Laura Maria Olimpi, Riccardo Ongaro, Maria Maddalena Palma, Angela Pasinato, Andrea Passarella, Pasquale Pazzola, Monica Perin, Danilo Perri, Silvana Rosa Pescosolido, Giovanni Petrazzuoli, Giuseppe Petrotto, Patrizia Picco, Ambrogina Pirola, Lorena Pisanello, Daniele Pittarello, Elena Porro, Antonino Puma, Maria Paola Puocci, Andrea Righetti, Rosaria Rizzari, Cristiano Rosafio, Paolo Rosas, Bruno Ruffato, Lucia Ruggieri, Annamaria Ruscitti, Annarita Russo, Pietro Salamone, Daniela Sambugaro, Luigi Saretta, Vittoria Sarno, Valentina Savio, Nico Maria Sciolla, Rossella Semenzato, Paolo Senesi, Carla Silvan, Giorgia Soldà, Valter Spanevello, Sabrina Spedale, Francesco Speranza, Sara Stefani, Francesco Storelli, Paolo Tambaro, Giacomo Toffol, Gabriele Tonelli, Silvia Tulone, Angelo Giuseppe Tummarello, Cristina Vallongo, Sergio Venditti, Maria Grazia Vitale, Concetta Volpe, Francescopaolo Volpe, Aldo Vozzi, Giulia Zanon, Maria Luisa Zuccolo.
+Appendix A Dataset
+A.1 Records distributions among hospitals and LHUs
+Figures A.1 and A.2 show the distribution of records across hospitals and LUHs, highlighting the number of positive bronchiolitis cases and the percentage of records from pediatric ER/Departments.
+Figure A.1: A. Distribution of bronchiolitis cases among hospitals. B. Distribution of bronchiolitis cases among LHUs.
+Figure A.2: A.
+Distribution of records among hospitals (only hospitals with >50>50 records).
+B.
+Distribution of records among LHUs (only LHUs with >50>50 cases).
+For each hospital/LHU, the blue bar indicates the total number of records, the light blue bar the records from pediatric ER/departments and the value on top of the bars is the % of pediatric records in the hospital/LHU.
+A.2 Discharge letters example
+An example of a discharge letter is reported below.
+This letter is a synthetic example similar in the type of content to those present in the analyzed dataset.
+We report it in both Italian and English.
+SERVIZIO SANITARIO NAZIONALE - REGIONE VENETO
+Azienda ULSS n.
+Presidio Ospedaliero di
+Unita Operativa di Pediatria Degenze
+Data:
+Egregio Collega,
+si dimette in data odierna: nato/a il
+Data e ora di ricovero:
+Diagnosi:
+INSUFFICIENZA RESPIRATORIA IN BRONCOSPASMO
+Decorso Clinico e Conclusioni:
+E.O. allingresso: condizioni generali sufficienti, Sat.O2 = 96%, FC = 154/m’. F.R.= 50/m’, P.A. = 100/60. Kg = 13,250, T = 36,3C
+Paziente pallido, polipnoica, al torace ridotta penetrazione daria, con ronchi e gemiti espiratori sparsi. Azione cardiaca valida e
+ritmica, addome piano, trattabile, non dolente ne dolorabile sia alla palpazione superficiale che profonda in tutti i quadranti, organi
+ipocondriaci nei limiti, peristalsi vivace. Faringe leggermente iperemico. Es. neurologico negativo. Non segni meningei.
+Sono stati eseguiti i seguenti esami, risultati negativi:
+Es. emocromocitometrico, azotemia, creatininemi, elettroliti, transaminasi, PCR, emogasanalisi.
+E stato riscontrato un modesto aumento della glicemia, dovuta allasomministrazione di cortisone.
+E stata iniziata terapia con O2 e con aerosol con broncovaleas 8 gtt + sol. fisiologica una somministrazione ogni 30 m per 3 volte,
+seguita da somministrazione dello stesso prodotto ( 8 gtt) 8 volte al di (ogni 3 h), con buona risposta da parte della sintomatologia
+respiratoria. Alla terapia aerosolica e stata associata terapia cortisonica per os con bentelan, 1,0 mg 2 volte al di’.
+i
+Programma alla dimissione:
+A domicilio si consiglia:
+- Aerosol con broncovaleas 6 gtt + sol. fisiol. 3 - 4 volte al di per 3 - 5 giorni.
+- Lavaggi nasali con sol. fisiol. sei volte al di per 3 - 5 giorni.
+- Bentelan 2 cp da 0,5 mg alla sera, a stomaco pieno, per 2 giorni.
+Consigli Clinici:
+Si consiglia inoltre controllo presso il pediatra curante nei prossimi giorni.
+A disposizione per ulteriori chiarimenti, inviamo cordiali saluti
+Cordiali saluti
+Dr.
+PED1D Pediatria Degenze
+INFORMAZIONE
+Gentile signore/ signora, desideriamo renderla partecipe che per il suo percorso di cura e stato stimato un impiego di risorse
+economiche da parte del Servizio Sanitario Regionale pari a Euro 1.945,69
+Paziente: ( ID: )
+Episodio N.: /01 ( Unita Operativa di PED1D Pediatria Degenze )
+stampato il Pag.1 of 1
+Referto firmato digitalmente.
+NATIONAL HEALTH SERVICE - VENETO REGION
+ULSS Company No.
+Hospital of
+Pediatric Unit Hospitalization
+Date:
+Dear Colleague,
+discharged today: born on
+Date and time of admission:
+Diagnosis:
+RESPIRATORY FAILURE IN BRONCHOSPASM
+Clinical Course and Conclusions:
+Upon admission: general condition sufficient, Sat.O2 = 96%, HR = 154/m’, $R$ = 50/m’, BP = 100/60. Weight = 13,250 kg, T = 36.3C
+Patient pale, tachypneic, with reduced air entry in the chest, with scattered wheezing and expiratory moaning. Valid and rhythmic
+cardiac action, flat abdomen, treatable, not painful or tender on both superficial and deep palpation in all quadrants, hypochondriac
+organs within limits, active peristalsis. Pharynx slightly hyperemic. Neurological exam negative. No meningeal signs.
+The following tests were performed, all negative:
+Blood count, urea, creatinine, electrolytes, transaminases, CRP, blood gas analysis.
+A slight increase in blood glucose was found, due to the administration of cortisone.
+Therapy was started with O2 and aerosol with broncovaleas 8 drops + saline solution, one administration every 30 minutes for 3 times,
+followed by administration of the same product (8 drops) 8 times a day (every 3 hours), with good response from respiratory symptoms.
+Aerosol therapy was combined with oral cortisone therapy with bentelan, 1.0 mg twice a day.
+i
+Discharge Plan:
+At home, it is advised:
+- Aerosol with broncovaleas 6 drops + saline solution 3 - 4 times a day for 3 - 5 days.
+- Nasal washes with saline solution six times a day for 3 - 5 days.
+- Bentelan 2 tablets of 0.5 mg in the evening, on a full stomach, for 2 days.
+Clinical Recommendations:
+It is also advised to check with the primary care pediatrician in the coming days.
+Available for further clarification, we send our best regards.
+Best regards,
+Dr.
+PED1D Pediatric Hospitalization
+INFORMATION
+Dear sir/madam, we wish to inform you that an estimated expenditure of resources by the Regional Health Service for your care is
+Euro 1,945.69.
+Patient: ( ID: )
+Episode No.: /01 ( Pediatric Unit of PED1D Hospitalization )
+printed on Page 1 of 1
+Digitally signed report.
+A.3 E3C Corpus
+Table A.1 reports the details about the composition of the E3C corpus.
+Table A.1: Number of documents in the E3C corpus for the different languages and types of sources
+[…]
+pubmed 9318 132 0 0 0
+journal 716 1637 798 504 201
+SPACCC corpus 0 0 0 1645 0
+other 0 24220 9675 0 1232
+Appendix B Pre-processing
+From diagnosis strings we remove punctuation, dates and times, and we trim new lines and spaces in excess.
+To remove headers and footers from the text of whole letters we report below the strings used as identifiers of sentences to be removed.
+regione
+azienda
+dipartimento*
+pronto soccorso pediatrico*
+u.o.c*
+UOC *
+operativa*
+accettazione*
+u.o*
+UO *
+uo.*
+U.L.S.S.*
+ULSS*
+ospedale*
+presidio ospedaliero*
+via*
+punto di primo intervento
+pediatria*
+verbale*
+pag.
+pag
+pagina
+nato*
+tess.san
+codice fiscale
+comune di nascita
+cap
+indirizzo
+cartella dea
+documento firmato digitalmente
+informazione ai sensi
+il medico dimettente
+gentile signor
+copia di documento firmato e conservato
+desideriamo renderLa partecipe
+modulo di pronto soccorso
+direttore*
+ai genitori
+al medico
+indirizzo
+residente
+residenza
+nome
+cognome
+firma
+consegnare al proprio pediatra
+lorario di alcune prestazioni
+verbale di pronto soccorso
+della cartella*
+modulo di
+numero di certificato
+firmatario
+il referto e\ conservato
+ID Documento
+gentile signore
+informazione
+dettagli paziente
+verbale n
+cartella DEA
+priorita*
+tel
+fax
+indirizzo
+residenza
+domicilio
+segreteria
+data e ora
+Those with * must be removed only when they are strictly at the beginning of the document. Otherwise, they might not necessarily indicate a header sentence.
+Figure B.1 reports the difference in results with and without this preprocessing.
+Figure B.1: AUC ($\pm 95\%$ C.I.) on gold labels with and without pre-processing
+Appendix C Diagnosis extraction
+To extract diagnosis strings, we initially consider the following keywords to be looked for at the beginning of a sentence
+diagnosi
+diagnosi di dimissione
+diagnosi alla dimissione
+diagnosi testuale
+We initially take as a diagnosis string the sentence that follows the keywords we identified among the ones above.
+It can be on the same line, if present, or on the following line.
+Then, this candidate diagnosis string is cleaned by removing the portion(s) related to the following expressions:
+decorso clinico
+consigli terapeutici
+consiglio
+controllo
+a domicilio
+[…]
+These keywords are followed by indications related to therapies to be followed after discharge or additional details of the hospital stay, which are not strictly part of the diagnosis string.
+In particular, if we identify these keywords, we remove the portion of the sentence that starts from them up to the next major punctuation sign (dot or semicolumn).
+The remaining portion, if any, is extracted as a diagnosis string.
+Table C.1 reports the most frequently extracted diagnosis sentences.
+Table C.1: Most common diagnosis sentences extracted from discharge letters
+(frequency >>100)
+Extracted Diagnosis [ITA] Extracted Diagnosis [ENG] N. Occ.
+virosi virosis 598
+079.99 infezioni virali,non specificate 079.99 viral infections,unspecified 351
+nato singolo,nato in ospedale senza menzione di taglio cesareo single-born child,born in hospital without mention of cesarean section 324
+gastroenterite gastroenteritis 323
+trauma cranico minore minor head injury 212
+febbre fever 190
+trauma cranico non commotivo non-concussive head injury 174
+dolore addominale abdominal pain 170
+780.6 febbre 780.6 fever 139
+orticaria hives 135
+broncospasmo bronchospasm 117
+519.11 broncospasmo acuto 519.11 acute bronchospasm 116
+febbre di recente insorgenza fever of recent onset 112
+gastroenterite acuta acute gastroenteritis 109
+otite otitis 106
+trauma cranico head injury 106
+009.1 colite,enterite e gastroenterite di presunta origine infettiva 009.1 colitis,enteritis and gastroenteritis of presumed infectious origin 104
+vomito vomit 104
+flogosi delle alte vie respiratorie upper respiratory tract inflammation 102
+bronchiolite bronchiolitis 102
+Appendix D Additional clustering results
+Figure D.1: Adjusted Rand Index (ARI) (A) and Normalized Mutual Information (NMI) (B) among clusters obtained with different algorithms (HDBSCAN, K-Means, Agglomerative Hierarchical Clustering), with different hyperparameters.
+For HDBSCAN mcs = minimum cluster size, ms=minimum number of samples.
+For K-Means and Agglomerative Hierarchical Clustering, kmin, kmed and kmax use as number of clusters the minimum, median or maximum number of clusters found by HDBSCAN, respectively.
+Figure D.2: AUCs ($\pm 95\%$ C.I.) obtained in 10-fold cross validation using weak labels derived by applying different clustering algorithms (HDBSCAN, K-Means, Agglomerative Hierarchical Clustering), with different hyperparameters.
+For HDBSCAN mcs = minimum cluster size, ms=minimum number of samples.
+For K-Means and Agglomerative Hierarchical Clustering, kmin, kmed and kmax use as number of clusters the minimum, median or maximum number of clusters found by HDBSCAN, respectively.
+To assess the robustness of the weak‐label generation process, we compared clusterings obtained using multiple HDBSCAN, k-means, and agglomerative hierarchical clustering configurations.
+As shown in Figure D.1, HDBSCAN solutions exhibited very high internal consistency across parameter settings (ARI and NMI often >0.7>0.7), indicating that the underlying structure captured by the algorithm is stable.
+In contrast, clusterings from k-means and agglomerative hierarchical clustering formed separate blocks with substantially lower agreement with HDBSCAN (ARI 0.2–0.4), suggesting that these methods partition the data according to different inductive biases.
+We then evaluated the downstream utility of the derived weak labels in a supervised classification task (Figure D.2).
+Weak labels obtained from HDBSCAN consistently yielded the highest AUC values, whereas those from k-means and Ward resulted in lower performance.
+Although HDBSCAN configurations displayed broadly similar performance, some parameter choices led to noticeably larger uncertainty in downstream AUC estimates.
+This suggests that larger minimum cluster sizes and intermediate values of min_samples provide more stable weak labels and should be preferred as a starting point.
+Another additional comparison that we report here is otbained by varying the embedding model.
+In particular, we experiment using the embeddings produced by the original Umberto model.
+This leads to 488 clusters at the first level and 119 at the second level.
+Tables D.1 and D.2 describe the twenty largest clusters identified with the Umberto-TPT and the original Umberto, respectively.
+Table D.3 reports the main first-level clusters related to respiratory infections and Table D.4 the corresponding second-level clusters, using the original Umberto model.
+Table D.1: Keyword-based representation of the twenty largest first-level diagnosis clusters obtained with Umberto-TPT
 Cluster size Cluster keywords [ITA] Cluster keywords [ENG]
@@
 53 contusione piede destro dito right foot toe contusion
-Table 2: Keyword-based representation of the first-level clusters related to respiratory infections, with the % of bronchiolitis cases in each cluster and the % to the total number of cases in the dataset
-Cluster # Cluster keywords [ITA] Cluster keywords [ENG]
-Within cluster
-bronch. %
+Table D.2: Keyword-based representation of the twenty largest first-level diagnosis clusters from the original Umberto representation
+Cluster size Cluster keywords [ITA] Cluster keywords ENG
+615 trauma cranico head injury
+498 nato taglio born section
+405 diagnosi reazione locale vaccino diagnosis specific vaccine reaction
+368 infezioni specificate specified infections
+333 ansia cellulite diarrea febbre frattura influenza vomito anxiety cellulite diarrhea fever fracture flu vomiting
+238 aeree flogosi vie flogosis airways
+232 congiuntivite specificata painful pronation
+226 flogosi vie flogosis airways
+206 appendicite congiuntivite enterite rinite appendicitis conjunctivitis enteritis rhinitis
+199 allergica orticaria pregressa previous allergic orticaria
+195 coxalgia otalgia coxalgia otalgia
+178 broncospasmo episodio bronchospasm episode
+178 broncospasmo laringospasmo bronchospasm laryngospasm
+170 d insetto puntura insect bite
+148 acuta gastroenterite acute gastroenteritis
+144 alte infezione respiratorie vie high airways infection
+140 enterite gastroenterite enteritis gastroenteritis
+136 faringotonsillite virale viral pharyngotonsillitis
+135 insorgenza recente recent onset
+133 acuta bronchiolite bronchite acute bronchiolitis bronchitis
+Table D.3: Keyword-based representation of the first-level clusters related to respiratory infections, with the % of bronchiolitis cases in each cluster and the % to the total number of cases in the dataset from the original Umberto representation
+Cluster # Cluster keywords [ITA] Cluster keywords [ENG] Within cluster bronch. %
 % of total bronch.
 covered
-1 bronchiolite lieve mild bronchiolitis 85.19% 6,87%
-2 acuta bronchiolite iniziale lieve acute mild initial bronchiolitis 85.14% 44,48%
-3 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 74.05% 11,94%
-4 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 70,00% 15,02%
-5 alimentazione bronchiolite difficolta’ feeding difficulties bronchiolitis 50.00% 2,09%
-6 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 12.7% 2,39%
-7 broncospasmo infezione respiratoria bronchospasm respiratory infection 6.67% 0,30%
-8 broncospasmo disidratazione lieve bronchospasm mild dehydration 1.17% 0,60%
-9 flogosi alte vie respiratorie febbre high airways flogosis fever $0$,00%
-10 broncospasmo infezione vie respiratorie bronchospasm airways infection $0$,00%
-11 acuto broncospasmo acute bronchospasm $0$,00%
-12 broncospasmo flogosi vie bronchospasm airways flogosis $0$,00%
-13 broncospasmo polmonite bronchospasm pneumonia $0$,00%
-14 broncospasmo corso ongoing bronchospasm $0$,00%
-15 broncopolmonite sinistra left bronchopneumonia $0$,00%
-16 febbricola febbrile infezione respiratoria virosi febrile fever respiratory infection virosis $0$,00%
-17 broncospasmo episodio bronchospasm episode $0$,00%
-18 infezione alte vie respiratorie high airways infection $0$,00%
-19 flogosi infezione alte vie respiratorie flogosis high airways infection $0$,00%
-20 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis $0$,00%
-21 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm $0$,00%
-22 broncospasmo otite bronchospasm otitis $0$,00%
-23 broncospasmo parainfettivo bronchospasm parainfective $0$,00%
-24 broncospasmo virosi bronchospasm virosis $0$,00%
-25 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory $0$,00%
-26 crisi epilessia infezione paziente polmonare respiratoria seizure epilepsy respiratory infection lung patient $0$,00%
-27 altre infezioni acute vie respiratorie other acute respiratory infection airways $0$,00%
-Table 3: Keyword-based representation of the second-level clusters related to respiratory infections, with the % of bronchiolitis cases in each cluster and the % to the total number of cases in the dataset
-2nd level cluster # 2nd level clusters keywords [ITA] 2nd level clusters keywords [ENG] 1st level cluster # 1st level clusters keywords [ITA] 1st level clusters keywords [ENG] Within cluster bronch.
-% % of total bronch.
-covered 1 bronchiolite lieve mild bronchiolitis 1 acuta bronchiolite iniziale lieve acute mild initial bronchiolitis 85.16 % 51.34 % 2 bronchiolite lieve mild bronchiolitis 2 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 3 acuta bronchiolite corso insufficienza respiratoria acute bronchiolitis undergoing respiratory failure 74.05 % 11.94 % 3 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 4 acuto broncospasmo corso febbre paziente acute bronchospasm ongoing fever patient 70.00 % 15.02 % 4 bronchiolite difficolta di alimentazione feeding difficulties bronchiolitis 5 bronchiolite difficolta di alimentazione feeding difficulties bronchiolitis 50.00 % 2.09 % 5 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 6 acuta infezione insufficienza respiratoria transitoria virosi acute respiratory infection transient virosis 12.7 % 2.39 % 6 broncospasmo infezione respiratoria bronchospasm respiratory infection 7 broncospasmo infezione vie respiratorie bronchospasm infection airways 2.9 % 0.30 % 10 broncospasmo infezione respiratoria bronchospasm respiratory infection 7 broncospasmo disidratazione lieve bronchospasm mild dehydration 8 broncospasmo disidratazione lieve bronchospasm mild dehydration 1.17 % 0.60 % 8 broncospasmo bronchospasm 11 acuto broncospasmo acute bronchospasm 0.00 % 0.00 % 12 broncospasmo flogosi vie bronchospasm airways flogosis 13 broncospasmo polmonite bronchospasm pneumonia 14 broncospasmo corso ongoing bronchospasm 22 broncospasmo otite bronchospasm otitis 23 broncospasmo parainfettivo bronchospasm parainfective 24 broncospasmo virosi bronchospasm virosis 17 broncospasmo episodio bronchospasm episode 9 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis 20 acuta broncopolmonite linfadenite acute bronchopneumonia lymphadenitis 0.00 % 0.00 % 10 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm 21 acuto broncospasmo tosse verosimilmente cough likely acute bronchospasm 0.00 % 0.00 % 11 infezione vie respiratorie airways infection 18 infezione alte vie respiratorie high airways infection 0.00 % 0.00 % 19 flogosi infezione alte vie respiratorie flogosis high airways infection 0.00 % 0.00 % 27 altre infezioni acute vie respiratorie other acute respiratory infection airways 0.00 % 0.00 % 26 crisi epilessia infezione paziente polmonare respiratoria seizure epilepsy respiratory infection lung patient 0.00 % 0.00 % 16 febbricola febbrile infezione respiratoria virosi febrile fever respiratory infection virosis 0.00 % 0.00 % 12 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory 25 infezione lattante pianto pregresso respiratoria infant infection infant cries previous respiratory 0.00 % 0.00 % 13 alte febbre flogosi respiratorie vie high fever flogosis airways 9 flogosi alte vie respiratorie febbre high airways flogosis fever 0.00 % 0.00 % 14 broncopolmonite sinistra left bronchopneumonia 15 broncopolmonite sinistra left bronchopneumonia 0.00 % 0.00 %
-3.4 Disease classification
-[…]
-[…]
-[…]
-When comparing the classification results on weak labels (W) and gold labels (G), we observe a reduction in precision, affecting both the F1-Score and the AUC.
-However, recall is higher on gold labels, at least for BERT-based models, whereas simpler models also showed a decrease in recall.
-The BERT-based models show minimal differences between settings with and without diagnosis string, indicating that the information needed to identify the diagnosis of bronchiolitis is present in the entire letter text, and these models can effectively recognize it.
-As expected, the fully supervised model demonstrated better overall performance, though differences are modest, particularly in recall, which is lower on average than our best weakly-supervised model.
-In Table 5, we assess the differences in the results of our best model when selecting different subsets of clusters for the weak labels.
-In particular, we exclude one of the 1st level clusters we selected at a time.
-Excluding clusters #5 or #1 does not affect the results, which is reasonable considering their smaller size.
-Excluding cluster #2 determines a drop in the performance and an increase in the variance, which is expected since it accounts for almost half of the bronchiolitis cases.
-Removing clusters #3 or #4 has a negative effect, but limited.
-These results confirm the solidity of our weak labels and the robustness of the model with respect to small changes in them.
-Table 6 compares metrics at the hospital level, revealing notable differences.
-Some hospitals, such as H-0 and H-3, are more represented in the dataset, which might explain their superior performance.
-However, hospitals with a similar number of records show considerable performance variability, suggesting that reporting practices in discharge letters might influence disease detection.
-A comparison between records from pediatric vs non-pediatric ER/Departments revealed significant differences (gold AUC $80.89\pm 6.37$ vs $71.07\pm 4.57$, see Supplementary Material E for details), but there are also hospitals with pediatric ER and low performances, such as H-4.
-Similar findings are reported in Supplementary Material E for the different LHUs.
-Table 4: 10-fold CV results for classification, comparing with rule-based baselines, fully supervised approaches and other classification modules in our pipeline.
-Std.
-Dev.
-in parenthesis.
-G = gold labels, W = weak labels.
-With diagnosis string
-Model
-P [%]
-R [%]
-F1 [%]
-AUC [%]
-P [%]
-R [%]
-F1 [%]
-AUC [%]
-Umberto-E3C 82.30 (8.94) 75.46 (6.25) 77.33 (4.06) 82.30 (4.99) 72.75 (7.50) 84.59 (6.48) 75.06 (3.72) 77.68 (5.73)
-Umberto 80.67 (5.99) 73.48 (7.12) 76.91 (5.97) 82.56 (4.79) 62.29 (6.24) 75.99 (4.69) 68.46 (2.77) 69.82 (4.54)
-LSTM 68.57 (6.48) 64.97 (11.96) 65.54 (5.66) 67.50 (3.99) 57.39 (11.37) 60.86 (13.67) 55.08 (8.07) 63.89 (5.28)
-Rule-based full text 70.18 87.05 77.70 - 65.57 68.76 67.13 -
-Rule-based diagnosis 73.88 86.38 79.64 - 74.61 67.42 70.83 -
-Umberto-E3C Supervised - - - - 80.30 (7.23) 81.52 (6.70) 80.91 (2.81) 86.45 (2.82)
-Without diagnosis string
-Model
-P [%]
-R [%]
-F1 [%]
-AUC [%]
-P [%]
-R [%]
-F1 [%]
-AUC [%]
-Umberto-E3C 84.34 (5.19) 73.04 (10.53) 75.15 (4.27) 84.91 (3.50) 70.15 (6.40) 79.00 (5.82) 74.28 (4.10) 77.56 (3.11)
-Umberto 80.01 (6.84) 68.68 (10.55) 73.90 (5.36) 80.81 (6.24) 61.41 (6.09) 76.69 (4.22) 68.18 (3.62) 70.49 (3.89)
-LSTM 64.98 (6.89) 56.21 (18.88) 58.80 (13.52) 63.00 (9.24) 50.14 (8.31) 63.97 (13.52) 53.03 (9.69) 59.02 (7.36)
-Rule-based full text 65.54 79.16 71.69 - 62.33 60.05 61.10 -
-Umberto-E3C Supervised - - - - 75.56 (5.08) 82.55 (9.14) 78.90 (4.54) 84.05 (4.26)
-Table 5: 10-fold CV results for classification, considering different groups of 1st level clusters for weak labels. Std. Dev. in parenthesis. G = gold labels, W = weak labels
-Clusters for weak labels
-(1st level)
-P [%]
-R [%]
-F1 [%]
-AUC [%]
-P [%]
-R [%]
-F1 [%]
-AUC [%]
-#1. #2. #3. #4 78.39 (7.04) 73.35 (6.21) 75.79 (4.65) 83.02 (4.16) 70.75 (6.48) 81.27 (6.99) 75.65 (4.29) 75.49 (5.57)
-#1. #2. #4. #5 73.69 (8.89) 83.20 (9.16) 78.16 (7.80) 80.71 (7.01) 68.42 (8.70) 75.09 (8.58) 71.60 (7.26) 72.67 (6.87)
-#2. #3. #4. #5 83.49 (7.87) 74.66 (6.88) 78.82 (4.26) 84.02 (4.43) 68.54 (6.30) 82.35 (4.94) 74.81 (3.82) 75.83 (5.22)
-#1. #3. #4. #5 93.73 (6.85) 78.11 (8.47) 85.19 (7.06) 90.91 (7.35) 57.29 (12.97) 59.03 (15.94) 58.15 (12.45) 68.90 (10.04)
-#1. #2. #3. #5 80.45 (8.90) 74.62 (8.57) 77.43 (5.76) 85.95 (5.29) 67.86 (6.36) 73.42 (8.64) 70.53 (5.33) 70.22 (3.10)
-Table 6: 10-fold CV results for classification, comparing different hospitals (only hospitals with more than 15 bronchiolitis cases). G = gold labels, W = weak labels.
+1 acuta bronchiolite lieve severa acute mild severe bronchiolitis 87.76 12.84
+2 acuta bronchiolite bronchite acute bronchiolitis bronchitis 72.93 28.96
+3 bronchiolite respiratorio virus bronchiolitis respiratory virus 66.66 2.99
+4 bronchiolite positiva vrs bronchiolitis rsv positive 44.44 1.19
+5 bronchiolite complicata corso ex gastroenterite paziente bronchiolitis complicated course former gastroenteritis patient 40.00 0.60
+6 bronchite broncospasmo lieve bronchitis mild bronchospasm 20.00 0.30
+7 bronchiolite broncospasmo lieve bronchiolitis mild bronchospasm 16.67 0.30
+8 allarme bronchiolite caratteristiche d epistassi insetto locale lussazione medicazione pregressa reazione sdr sn vrs bronchiolitis alarm features epistaxis local insect dislocation dressing previous reaction sdr sn rsv 16.67 0.30
+9 acuta insufficienza respiratoria acute respiratory failure 14.04 2.39
+10 broncospasmo respiratoria bronchospasm respiratory 12.50 2.99
+11 broncospasmo laringospasmo bronchospasm laryngospasm 1.12 0.60
+12 broncospasmo episodio bronchospasm episode 0.56 0.30
+13 febbrile respiratoria respiratory fever 0.0 0.0
+14 febbre recente respiratorie vie recent fever airways 0.0 0.0
+15 corso infezione respiratoria ongoing respiratory infection 0.0 0.0
+16 broncospasmo vie bronchospasm airways 0.0 0.0
+17 broncospasmo paziente ricorrente bronchospasm recurrent patient 0.0 0.0
+18 broncospasmo parainfettivo parainfective bronchospasm 0.0 0.0
+19 broncospasmo noto known bronchospasm 0.0 0.0
+20 broncospasmo infettivo parainfettivo parainfective infectious bronchospasm 0.0 0.0
+21 broncospasmo corso virosi bronchospasm ongoing virosis 0.0 0.0
+22 broncospasmo corso respiratorie vie bronchospasm ongoing airways 0.0 0.0
+23 broncospasmo corso infezione respiratoria bronchospasm ongoing respiratory infection 0.0 0.0
+24 broncospasmo corso flogosi vie bronchospasm ongoing flogosis airways 0.0 0.0
+25 broncospasmo corso flogosi lieve respiratoria bronchospasm ongoing mild flososis respiratory 0.0 0.0
+26 broncopolmonite destra iniziale initial right bronchopneumonia 0.0 0.0
+27 batterica broncopolmonite esterna infettiva otite polmonite specificata bacterial bronchopneumonia external infectious otitis pneumonia specified 0.0 0.0
+28 alte vie respiratorie infezione high airways infection 0.0 0.0
+29 alte vie respiratorie flogosi high airways flogosis 0.0 0.0
+30 corso alte vie respiratorie flogosi ongoing high airways flogosis 0.0 0.0
+31 corso alte vie respiratorie virosi febbre ongoing high airways virosis fever 0.0 0.0
+32 acuto broncospasmo lieve moderato acute mild moderate bronchospasm 0.0 0.0
+33 acute infezioni vie respiratorie superiori acute high airways infection 0.0 0.0
+34 acuta alte vie respiratorie rinite virosi acute high airways virosis rhinitis 0.0 0.0
+Table D.4: Keyword-based representation of the second level clusters related to respiratory infections from the original Umberto representation
+2nd level cluster # 2nd level clusters keywords [ITA] 2nd level clusters keywords [ENG] 1st level cluster # 1st level clusters keywords [ITA] 1st level clusters keywords [ENG]
+1 acuta corso infezione insufficienza respiratoria acute bronchiolitis undergoing respiratory failure # 9 acuta insufficienza respiratoria acute respiratory failure
+# 15 corso infezione respiratoria ongoing respiratory infection
+2 broncospasmo episodio noto vie known bronchospasm episode airways # 16 broncospasmo vie bronchospasm airways
+# 19 broncospasmo noto known bronchospasm
+# 12 broncospasmo episodio episode bronchospasm
+3 broncospasmo corso paziente ricorrente virosi bronchospasm ongoing recurrent patient virosis # 21 broncospasmo corso virosi bronchospasm ongoing virosis
+# 17 broncospasmo paziente ricorrente bronchospasm recurrent patient
+4 broncospasmo lieve bronchospasm mild # 6 bronchiolite broncospasmo lieve bronchiolitis bronchospasm mild
+# 7 bronchiolite broncospasmo lieve bronchiolitis bronchospasm mild
+# 32 acuto broncospasmo lieve moderato acute bronchospasm mild moderate
+5 acuta bronchiolite acute bronchiolitis # 2 acuta bronchiolite bronchite acute bronchiolitis bronchitis
+# 1 acuta bronchiolite lieve severa acute bronchiolitis mild severe
+6 respiratorie vie airways # 28 alte infezione respiratorie vie high airways infection
+# 31 alte vie respiratorie corso febbre virosi high airways ongoing fever virosis
+# 33 acute infezioni vie respiratorie superiori acute high airways infection
+# 34 acuta alte respiratorie rinite vie virosi acute high airways rhinitis virosis
+# 14 febbre recente respiratorie vie recent fever airways
+8 broncospasmo respiratoria bronchospasm respiratory # 23 broncospasmo corso infezione respiratoria bronchospasm ongoing respiratory infection
+# 25 broncospasmo corso flogosi lieve respiratoria bronchospasm ongoing flogosis mild respiratory
+# 10 broncospasmo respiratoria bronchospasm respiratory
+9 broncospasmo corso vie bronchospasm ongoig airways # 24 broncospasmo corso flogosi vie bronchospasm ongoing flogosis
+# 22 broncospasmo corso respiratorie vie bronchospasm ongoing airways
+10 alte flogosi respiratorie vie flogosis high airways # 30 corso flogosi alte vie respiratorie flogosis high airways ongoing
+# 29 flogosi alte vie respiratorie flogosis high airways
+11 bronchiolite respiratorio virus bronchiolitis respiratory virus # 3 bronchiolite respiratorio virus bronchiolitis respiratory virus
+12 bronchiolite positiva vrs bronchiolitis rsv positive # 4 bronchiolite positiva vrs bronchiolitis rsv positive
+13 bronchiolite complicata corso ex gastroenterite paziente bronchiolitis complicated course former gastroenteritis patient # 5 bronchiolite complicata corso ex gastroenterite paziente bronchiolitis complicated course former gastroenteritis patient
+14 allarme bronchiolite caratteristiche d epistassi insetto locale lussazione medicazione pregressa reazione sdn sn vrs bronchiolitis alarm features epistaxis local insect dislocation dressing previous reaction sdn sn rsv # 8 allarme bronchiolite caratteristiche d epistassi insetto locale lussazione medicazione pregressa reazione sdn sn vrs bronchiolitis alarm features epistaxis local insect dislocation dressing previous reaction sdn sn rsv
+15 febbrile respiratoria respiratory fever # 13 febbrile respiratoria respiratory fever
+16 broncopolmonite destra iniziale initial right bronchopneumonia # 26 broncopolmonite destra iniziale initial right bronchopneumonia
+17 batterica broncopolmonite esterna infettiva otite polmonite specificata bacterial bronchopneumonia external infectious otitis pneumonia specified # 27 batterica broncopolmonite esterna infettiva otite polmonite specificata bacterial bronchopneumonia external infectious otitis pneumonia specified
+Appendix E Additional classification results
+In this section, we report additional classification results to assess the effect and the differences among hospitals and LHUs, in Tables E.1 and E.2, respectively.
+Since the number of cases for each hospital and LHU is limited, we consider those with a minimum number of bronchiolitis cases (15) and measure the results using a Leave One Group Out (LOGO) approach.
+For each hospital/LHU to be evaluated, we train the model on the data from the other hospitals/LHUs, and we measure the performance on the data from the hospital/LHU of interest.
+In Table E.3 we also compare results between the groups of pediatric vs non-pediatric ER/Department admissions.
+In this case, we adopt the same 10-fold stratified cross-validation used to compare different classification models, dividing the test set between pediatric and non-pediatric ERs/Departments at each iteration and computing the metrics on these two groups separately.
+We identified the records related to pediatric ERs/Departments by looking at the presence of specific keywords (pediatria, pediatrico, pediatrica) in the header of the letters.
+This marked 10’340 records (31.17%) as related to pediatric ERs/Departments.
+Table E.1: 10-fold CV results for classification, comparing different hospitals (only hospitals with more than 15 bronchiolitis cases). G = gold labels, W = weak labels.
 Hospital Dataset %
 Pediatric ER
+(% records)
 Bronchiolitis
+(% records)
 P [%]
@@
 H-6 2.59 9.40 2.09 100.00 31.25 47.62 65.63 100.00 33.33 50.00 70.48
-4 Discussion
-This study introduces a novel pipeline that shows how weakly-supervised learning can be effectively conducted in the clinical natural language processing domain.
-This is the first study to address the problem of diagnosis identification from discharge letters in the Italian language, and while we validated the method using a case study on bronchiolitis, the pipeline is designed to be applicable to any disease.
-Our case study shows excellent results, outperforming existing unsupervised approaches, and underscores the potential for broader application in clinical settings.
-We observed minimal performance differences between the metrics computed on weak and gold labels (1-2 % of F1-Score, 5-7 % of AUC), confirming that the first step of our pipeline (SAL) effectively identifies relevant cases.
-Despite the inevitable noise that the weak labels introduce, this is limited enough to allow the model to perform well when evaluated against gold labels.
-Additionally, the variations between settings where we retained the diagnosis strings and those where we removed them are negligible.
-This indicates that the model is effectively learning from the entire text of the discharge letters, making it particularly advantageous in situations where automatic extraction of diagnosis strings is not possible.
-Our manual exploration of some weak labels further revealed that diagnosis strings do not always mention the disease responsible for hospitalization.
-In some cases, they only refer to symptoms, and in cases of multiple diseases, not all may be reported.
-These findings highlight the necessity for a classification model based on the complete text of discharge letters, as the one we developed.
-In contrast, rule-based approaches exhibited poor performance, mainly due to the fact that diseases like bronchiolitis are not always explicitly mentioned by name in the text.
-Even when mentioned, they might refer to discarded hypotheses or contacts with affected individuals, leading to false positives.
-While a more sophisticated rule-based system could address some of these issues, such approaches lack the scalability and generalizability of our pipeline.
-Conversely, our pipeline can be easily adapted to other diseases, which makes it a versatile tool for clinical NLP.
-Our fine-tuned version of the Umberto model showed the best results both in the classification and the clustering modules, demonstrating a better understanding of the Italian medical domain when compared with the general-domain version.
-Compared to a fully supervised method, our weakly-supervised pipeline maintained competitive performance, with only a limited and acceptable loss ($\sim 5\%$ of F1-Score, $\sim 6\%$ of AUC).
-This is particularly notable given the significant time and effort required for manual annotation in fully supervised setups.
-The pipeline has also shown robustness to the cluster selection for the weak labels.
-These results underscore the potential of weakly-supervised approaches in clinical natural language processing, a domain where the scarcity of labelled data is the most relevant challenge.
-While our results are promising, we must acknowledge some limitations in our study that could be addressed in future work.
-The absence of a gold standard in labelling is a common challenge in clinical NLP, which underscores the value of weakly-supervised methods like ours.
-Our gold labels for bronchiolitis were manually assigned by individual physicians, making it impossible to verify inter-annotator agreement and increasing the risk of errors.
-The multi-centric nature of our dataset, though a strength in its diversity, also presents an imbalance in the number of records from different hospitals and Local Health Units.
-Our analysis highlighted relevant performance differences among hospitals and LHUs, warranting further investigation.
-Testing the pipeline on a more balanced dataset would strengthen conclusions about these performance differences.
-Moreover, this could facilitate the development of a hierarchical transformer-based classification model that considers hospital/LHU context in the input, potentially improving model performance.
-We also observed relevant differences between pediatric vs.
-non-pediatric ERs/Departments.
-Regarding this, we need to consider that pediatric admissions were identified using specific keywords.
-This method could miss some pediatric cases, although such instances should be limited.
-Future work could use data that directly includes this information as a structured field, ensuring its accuracy and completeness.
-Although our case study focused on bronchiolitis, the diagnostic clusters we identified appear capable of recognizing a broad range of diseases.
-Validating the pipeline on other diseases would further demonstrate its robustness and generalizability.
-Developing a multi-disease model, particularly for conditions that share similar symptoms, could provide broader applicability and enhance the model’s utility.
-Tailored rule-based models often fail in such scenarios, making a machine learning-based approach more effective.
-Another important direction for future research is enhancing the interpretability of our model.
-Identifying the portions of the discharge letter that are most relevant to the model’s output would help clinicians understand and trust the model’s decisions.
-Techniques related to the area of explainable AI could be employed to enhance interpretability, ensuring the model’s reasoning process is transparent and understandable.
-This would also help mitigate biases that could potentially be present in the training data, further enhancing the model’s reliability.
-5 Conclusions
-We have developed and validated the first weakly-supervised pipeline to analyze discharge letters in the Italian language, with a case study focusing on bronchiolitis.
-The pipeline achieved excellent results, outperforming rule-based approaches.
-Our fine-tuned BERT-based model showed superior performance in both the clustering and the classification module compared to a general-domain BERT model.
-We also assessed significant differences between hospitals/LHUs and pediatric vs.
-non-pediatric ERs, which suggest further investigations.
-However, while the case study showcased the method’s potential, the pipeline itself is designed to be adaptable to any disease, making it a valuable tool for researchers interested in extracting diagnoses from unstructured data and for epidemiological monitoring of diseases.
-Despite a small performance gap compared to fully supervised approaches, the weakly-supervised pipeline addresses real-world constraints, particularly the scarcity of labelled data in clinical NLP.
-Transformer-based models showed strong performance even without gold-labelled data, underscoring the viability of weakly-supervised methods in overcoming data limitations in the medical domain.
-Declaration of competing interests
-The authors declare that they have no competing interests.
-Data statement
-Data used in this study are collected and accessed through Pedianet, an Italian network of paediatricians.
-Data anonymization is performed in accordance with Italian regulations.
-Each anonymized record is assigned a unique numerical identifier.
-The data is then transmitted monthly to a centralized database located in Padua, Italy, for validation purposes.
-Inclusion in the Pedianet database is voluntary and parents/legal guardians must provide consent for their children’s anonymized data to be used for research purposes.
-Ethical approval of the study and access to the database were granted by the Internal Scientific Committee of So.Se.Te.
-Srl, the legal owner of Pedianet.
-Data used in this study are not publicly available due to their confidential nature. The code developed for their analysis is available by the authors upon request.
-CRediT authorship contribution statement
-Vittorio Torri: Methodology, Software, Investigation, Writing - Original Draft Elisa Barbieri: Conceptualization, Data Curation, Writing - Review & Editing Anna Cantarutti: Conceptualization, Data Curation, Writing - Review & Editing Carlo Giaquinto: Conceptualization, Supervision Francesca Ieva: Conceptualization, Methodology, Writing - Review & Editing, Supervision
-Acknowledgments
-The present research has been partially supported by MUR, grant Dipartimento di Eccellenza 2023-2027.
-Elisa Barbieri declares that her position is funded within the project “INF-ACT - One Health Basic and translational Research Actions addressing Unmet Needs on Emerging Infectious Diseases”, ID MUR PE_00000007, under the National Recovery and Resilience Plan (NRRP), Mission 4, Component 2 – CUP C93C22005170007.
-The authors thank all the family paediatricians collaborating in Pedianet: Eva Alfieri, Michela Alfiero Bordigato, Angelo Alongi, Biagio Amoroso, Rosaria Ancarola, Barbara Andreola, Giampaolo Anese, Roberta Angelini, Maria Grazia Apostolo, Giovanna Argo, Giovanni Avarello, Lucia Azzoni, Maria Carolina Barbazza, Patrizia Barbieri, Gabriele Belluzzi, Eleonora Benetti, Filippo Biasci, Franca Boe, Stefano Bollettini, Francesco Bonaiuto, Anna Maria Bontempelli, Matteo Bonza, Sara Bozzetto, Andrea Bruna, Ivana Brusaterra, Massimo Caccini, Laura Calì, Sonia Camposilvan, Laura Cantalupi, Luigi Cantarutti, Chiara Cardarelli, Giovanna Carli, Sylvia Carnazza, Massimo Castaldo, Stefano Castelli, Monica Cavedagni, Giuseppe Egidio Cera, Chiara Chillemi, Francesca Cichello, Giuseppe Cicione, Carla Ciscato, Mariangela Clerici Schoeller, Samuele Cocchiola, Giuseppe Collacciani, Valeria Conte, Roberta Corro’, Rosaria Costagliola, Nicola Costanzo, Sandra Cozzani, Giancarlo Cuboni, Giorgia Curia, Caterina D’alia, Vito Francesco D’Amanti, Antonio D’Avino, Roberto De Clara, Lorenzo De Giovanni, Annamaria De Marchi, Gigliola Del Ponte, Tiziana Di Giampietro, Giuseppe Di Mauro, Giuseppe Di Santo, Piero Di Saverio, Mattea Dieli, Marco Dolci, Mattia Doria, Dania El Mazloum, Maria Carmen Fadda, Pietro Falco, Mario Fama, Marco Faraci, Maria Immacolata Farina, Tania Favilli, Mariagrazia Federico, Michele Felice, Maurizio Ferraiuolo, Michele Ferretti, Mauro Gabriele Ferretti, Paolo Forcina, Patrizia Foti, Luisa Freo, Ezio Frison, Fabrizio Fusco, Giovanni Gallo, Roberto Gallo, Andrea Galvagno, Alberta Gentili, Pierfrancesco Gentilucci, Giuliana Giampaolo, Francesco Gianfredi, Isabella Giuseppin, Laura Gnesi, Costantino Gobbi, Renza Granzon, Mauro Grelloni, Mirco Grugnetti, Antonina Isca, Urania Elisabetta Lagrasta, Maria Rosaria Letta, Giuseppe Lietti, Cinzia Lista, Ricciardo Lucantonio, Francesco Luise, Enrico Marano, Francesca Marine, Lorenzo Mariniello, Gabriella Marostica, Sergio Masotti, Stefano Meneghetti, Massimo Milani, Stella Vittoria Milone, Donatella Moggia, Angela Maria Monteleone, Pierangela Mussinu, Anna Naccari, Immacolata Naso, Flavia Nicoloso, Cristina Novarini, Laura Maria Olimpi, Riccardo Ongaro, Maria Maddalena Palma, Angela Pasinato, Andrea Passarella, Pasquale Pazzola, Monica Perin, Danilo Perri, Silvana Rosa Pescosolido, Giovanni Petrazzuoli, Giuseppe Petrotto, Patrizia Picco, Ambrogina Pirola, Lorena Pisanello, Daniele Pittarello, Elena Porro, Antonino Puma, Maria Paola Puocci, Andrea Righetti, Rosaria Rizzari, Cristiano Rosafio, Paolo Rosas, Bruno Ruffato, Lucia Ruggieri, Annamaria Ruscitti, Annarita Russo, Pietro Salamone, Daniela Sambugaro, Luigi Saretta, Vittoria Sarno, Valentina Savio, Nico Maria Sciolla, Rossella Semenzato, Paolo Senesi, Carla Silvan, Giorgia Soldà, Valter Spanevello, Sabrina Spedale, Francesco Speranza, Sara Stefani, Francesco Storelli, Paolo Tambaro, Giacomo Toffol, Gabriele Tonelli, Silvia Tulone, Angelo Giuseppe Tummarello, Cristina Vallongo, Sergio Venditti, Maria Grazia Vitale, Concetta Volpe, Francescopaolo Volpe, Aldo Vozzi, Giulia Zanon, Maria Luisa Zuccolo.
-Declaration of generative AI and AI-assisted technologies in the writing process
-During the preparation of this work, the authors used ChatGPT to improve the readability and clarity of some sentences.
-After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
+Table E.2: 10-fold CV results for classification, comparing different LHUs (only LHUs with more than 15 cases). Std. Dev. in parenthesis. G = gold labels, W = weak labels.
+LHU
+P [%]
+R [%]
+F1 [%]
+AUC [%]
+P [%]
+R [%]
+F1 [%]
+AUC [%]
+LHU-3 75.56 65.38 70.10 78.81 69.70 53.49 60.53 73.94
+LHU-5 83.33 40.00 54.05 65.77 45.83 64.71 53.66 58.25
+LHU-2 94.12 64.00 76.19 88.89 51.61 94.12 66.67 59.95
+LHU-6 64.29 81.81 72.00 91.17 51.67 71.43 59.96 83.52
+LHU-7 86.27 83.02 84.62 87.06 82.50 76.74 79.52 86.22
+LHU-0 77.46 78.57 78.01 90.40 81.40 63.64 71.43 86.56
+Table E.3: 10-fold CV results for classification, comparing records from pediatric ERs/Departments with records from non-pediatric ERs/Departments.
+Std.
+Dev.
+in parenthesis.
+G = gold labels, W = weak labels.
+Pediatrics ER
+P [%]
+R [%]
+F1 [%]
+AUC [%]
+P [%]
+R [%]
+F1 [%]
+AUC [%]
+Yes 87.79 (7.12) 83.31 (4.30) 85.13 (6.56) 86.04 (5.29) 75.89 (9.85) 88.57 (8.23) 81.74 (7.71) 80.89 (6.37)
+No 72.47 (7.28) 80.52 (7.87) 76.08 (5.64) 75.93 (4.04) 60.29 (8.29) 85.21 (6.11) 69.37 (4.36) 71.07 (4.57)